### v1

# An approach to non-equilibrium statistical physics using variational Bayesian inference

Abstract
We discuss an approach to mathematically modelling systems made of objects that are coupled together, using generative models of the dependence relationships between states (or trajectories) of the things comprising such systems. This broad class includes open or non-equilibrium systems and is especially relevant to self-organising systems. The ensuing variational free energy principle (FEP) has certain advantages over using random dynamical systems explicitly, notably, by being more tractable and offering a parsimonious explanation of why the joint system evolves in the way that it does, based on the properties of the coupling between system components. Using the FEP allows us to model the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics. In short, we argue that using generative models to represent and track relations among subsystems leads us to a particular statistical theory of interacting systems. Conversely, this theory enables us to construct nested models that respect the known relations among subsystems. We point out that the fact that a physical object conforms to the FEP does not necessarily imply that this object performs inference in the literal sense; rather, it is a useful explanatory fiction which replaces the ‘explicit’ dynamics of the object with an ‘implicit’ flow on free energy gradients—a fiction that may or may not be entertained by the object itself.
Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna and the Laufer Centre for Physical and Quantitative Biology, and the informal discussions that followed.
1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [2, 3], describes how a physical object reflects the properties of its environment and vice versa. Bayesian mechanics under the free energy principle (FEP) [4, 5] aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference), resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled.
In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the FEP—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the FEP to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling. We identify a particular issue, which the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way.
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that the FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [6, 7]. Some have referenced this fallacy in relation to the FEP [8, 9]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [10, 11, 12, 3] for related critical discussions of such claims.)
According to this line of thinking, using Bayesian mechanics to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific FEP-theoretic or Bayesian mechanical model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [10, 11]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of anything whatsoever that exists in the physical world.
Here, we discuss the use of the FEP in statistics and statistical physics, and what advantages it offers; we then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished. We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy. This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful. We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object. That is to say, a model of an object can always be written as modelling or inferring the structure of its environment. To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [15]; see also [13, 16, 3]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [17], which captures conditional independence relations amongst subsets of the system [2, 18]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is (a form of) Bayesian inference. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
Accounts such as [10] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
We first present a conceptual, and then technical, introduction to the FEP and Bayesian mechanics—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other.
We propose that the FEP and Bayesian mechanics provide a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads, in a Wittgensteinian manner, to constraints on all possible maps, or models of physical things. This is meant in terms of logically necessary preconditions for sense-making, and thus the parsing of sensory streams on which cognition is contingent [15]. In particular, these constraints arise from what it means to be a map or model of a physical process at all.
We remark that we will not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]). Models built using the FEP are still scientific models, and the properties of FEP theoretic model also run the risk of reification. Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
2 The free energy principle
2.1 Conceptually
In this subsection we will provide an up-to-date overview of the concepts behind the FEP. One can also see [3] for a discussion of the main ideas. See the following subsection for more mathematical details.
Conceptually, the FEP starts from a partitions an overall “system” into three component sets of states: the internal states $\mu$ of a ‘thing’ or ‘particle’, an environment (which possesses external states $\eta$, relative to the thing in question), and an interface separating but coupling the two (the so-called Markov blanket, $b$), which mediates the influence of $\eta$ on $\mu$ and vice-versa. The states of the system will be tuples $x\coloneqq(\eta,b,\mu)$, where $\eta$ and $\mu$ are states of two subsystems and $b$ is some locus of states such that $p(\eta\mid\mu,b)=p(\eta\mid b)$ and vice versa. “Particular states” are the internal states of each thing and their respective Markov blanket. Of note, the particular partition is akin to choosing a set of coarse-grained variables at a superordinate scale, associating each state or path at the subordinate scale to a particle at the superordinate scale (as either an internal or a blanket state), in terms of the relations of conditional independence that obtain between states or paths [20, 2].
More precisely, we will consider systems of interacting objects, such as an open random dynamical system coupled to its environment. Fixing a final time $\tau$, the joint system will be denoted by the time-indexed family of random variables $\{X\}_{t:\tau}$ taking values $x\in\mathcal{X}$ at each $t$, with a specific realisation of the system (meaning, a sample path or trajectory beginning at some fixed $X_{0}=x_{0}$) being the sequence $\gamma(x_{0})=\{x_{0},x_{1},\ldots,x_{\tau}\}$. We will define $\gamma(x_{0})\mapsto\gamma_{t}(x_{0})$ as the projection to the state at time $t$. Then $B_{t}$ is what we have called the Markov blanket in $\{X\}_{t:\tau}$ at time $t$, separating it into statistically distinct subsystems (which one can then argue do not mix into each other [18]). We will repeat this notation for
$\gamma_{t}(\eta_{0})=N_{t}=\eta\in\mathcal{N}$
$\gamma_{t}(b_{0})=B_{t}=b\in\mathcal{B}$
$\gamma_{t}(\mu_{0})=M_{t}=\mu\in\mathcal{M}$
as well as for conditional trajectories such as
$\gamma_{t}(\eta_{0}\mid B_{t}=b)=\eta_{b}\in\mathcal{N}\crossproduct\{b\}.$
Although we will not go into the theory of random dynamical systems, when equipped with a time evolution map acting on the sample space, these are all the ingredients necessary to define a random dynamical system [21], and this is implicit in our discussion; for details see the account in [14]. When $\{M\}_{t:\tau}$ is an open system coupled to an environment with states $\eta$, a value $\mu$ is called an internal state, and the subsystem $\{M\}_{t:\tau}$ is variously called an object, a thing, or a particle in the literature. This is the convention we will stick to.
We can guarantee that the dynamics of a system can be described as a flow on surprisal, and thereby, cast as implementing a form of inference, whenever there exists a “particular partition” of that system. At the core of the FEP is a mathematical statement about the behaviour of such systems. For a comprehensive overview, see [3]. In particular, the FEP allows us relate the dynamics of states or paths that are internal to the object, $\mu$, to the dynamics of external states, conditioned on boundary or blanket states $p(\eta\mid b)$—that is, understanding what goes on inside a subsystem by understanding what it is coupled to, as ‘seen’ via blanket states [5, 22].
The logic of the FEP can be summarised as follows. If there exists a particular partition of a system, then there exists a coupling between internal and external states or paths, given blanket states or paths; then, since the internal dynamics of a given object are coupled to other (external) subsets of the system—in such a way that they can be read as parameters for probability densities over the external subset—the FEP describes these interacting systems using variational inference. This is a core aspect of the FEP that distinguishes it from a straightforward application of density dynamics to the internal states per se.
This is where the interpretation of dynamics as a form of inference comes into play. Indeed, one way to state the FEP is:
Claim 1.
If a random dynamical system (i) is subject to a specific kind of sparse coupling or boundary condition, such that it possesses a Markov blanket, and (ii) has a subsystem $\{M\}_{t:\tau}$ under that blanket with characteristic states (in the sense that it occupies a pullback attractor), then $\{M\}_{t:\tau}$ can always be described as if it minimises a variational free energy functional of densities over states or trajectories of $\{N\}_{t:\tau}$ [14, 22].
Another statement one can make is that Bayesian mechanics formalises the truism that, on average, an object will do whatever it is that it does on average, given the sort of object it is. Behind this tautological statement, we find a formal apparatus that enables us to model a variety of dynamical systems in a parsimonious manner. We hope this will allow us to take some first steps towards a mechanistic description of how self-organisation manifests in physical systems with such pullback attractors [3].
Bayesian mechanics explains why every physical thing looks as if it were performing inference over the causes of its perturbations, forming (Bayes) optimal beliefs about those causes [22]. Bayesian mechanics begins from the minimisation of self-information (also called surprisal) or free energy, and allows us to write down mechanical theories (i.e., laws of motion) yielding the dynamics of systems on an attractor [19]. Surprisal is simply the negative log-probability of an observed outcome. To say that an object minimises (or follows a path of least) surprisal is to say that an object evolves towards the outcomes which are most probable, given the sort of thing that it is (meaning, its characteristic states of being, or characteristic trajectories through its state space).
Interestingly, this means that surprisal is suited to act as a kind of ‘ontological potential’ for the object, in the sense that the surprisal is equivalent to a constraint on what states and outcomes are accessible to the object; the trajectories of a physical system are random curves through the gradient of surprisal over the state space, which satisfy an equation of motion derived from the surprisal itself [4, 23]. As has been asserted in recent literature [19], one reading of the FEP would hold that the crux of the ubiquity of the FEP is that everything which is coupled to another thing holds a model of that thing to which it is coupled in the following, trivial sense: in virtue of this coupling, one subsystem estimates the parameters of the probability density over states (or trajectories) of the other subsystems (and vice versa). Using this property, they can be written as performing inference in the sense of variational Bayesian inference.
This coupling is sometimes discussed in terms of the ‘probabilistic beliefs’ that one object holds about its environment. Given that the term ‘belief’ is quite theoretically loaded (especially in philosophy), we ought perhaps to say ‘estimator,’ instead. Once more, our contention is that anything which estimates the parameters of a density is engaged in a form of Bayesian inference, by mathematical definition; i.e., it is mathematically indisputable that systems that estimate statistics represent, or encode, or otherwise hold probabilistic beliefs—whether the practitioner likes the word ‘beliefs’ or not, and irrespective of what commitments one has with respect to notions such as encoding and representation (e.g., [24]).
Generative models and variational methods allow us to licence these claims [25, 26]. Generative models are joint probability densities over causes and consequences which, here, are used to capture the dependencies between the variables constituting $\{X\}_{t:\tau}$, meaning the way that changes in some variables cause changes in others. Writing down a generative model for some real object in the world (the metaphorical territory) is building a metaphorical map of that territory (a formalisation or codification of the relations between the things in that territory).
Variational techniques comprise a family of inference techniques which allow a modeller to approximate some (usually intractable) posterior probability density under a (usually specified) generative model. In variational inference, we write down a parameterised probability density function called a variational density, which is used as an approximation to the posterior probability density. The quantity called variational free energy scores the difference between the variational density and the posterior over external states—here, a conditional density, given a particular blanket state. Hence, when some notion of the difference between them is zero, variational free energy can be used as a proxy for surprise, and we can approximate the posterior density; which is a form of parameterised Bayesian inference.
Note that any given object cannot change the generative model associated to that object. This follows because the object can only change the variational density encoded by its internal states. However, to explain these changes—under a variational principle of least action—one needs to specify the generative model under which the posterior is approximated. This means the generative model explains the dynamics of something that is observed in a particular way. This is usually expressed by saying that the generative model is entailed by the thing in question. That an object entails a generative model is consistent with statements like “an object is a model of its environment”: e.g., in the sense of the good regulator theorem [19]). This follows because the generative model just is a probabilistic description of the joint states an object and its environment occupy.
2.2 Mathematically
With that conceptual overview being stated, we will now provide some further mathematical details. Similar (contemporary) reviews can be found in [4, 5, 19, 27]. We will argue that coupled random dynamical systems partitioned under a Markov blanket naturally have the form of variational Bayesian inference problems, in the sense that we obtain the values of parameters from the minimisation of free energy or surprisal. This is what licences the claim that minimising free energy implements a form of inference, in which case the object models its environment. Ultimately this simply means that, in virtue of being coupled via blanket states to its environment, the object reflects the structure of, and data about (i.e., models), its environment—or, what it ‘believes’ that its environment is like (in the sense of being a statistical estimator, as defined above).
Appealing to the Laplace approximation (that if a density has a unique maximum, in a neighbourhood of the most likely state, then the density is approximately Gaussian with that mean), we will suppose all random variables and all trajectories are jointly Gaussian. We will call the conditionally expected state—that is, the expected state given some fixed blanket state—the conditional mode $\hat{\mu}_{b}$ or $\hat{\eta}_{b}$. Let $\mathcal{L}(\mu,b)=-\log p(\mu,b)$ be the surprisal of internal and blanket states, such that $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$ and thus satisfies $\nabla_{\mu}\mathcal{L}(\mu,b)=0$. By the Freidlin–Wentzell theorem, we can argue the same is true of paths (see [19]). In that setting, the interpretation of surprisal as the fluctuations of a random dynamical system away from its expected path plays a crucial role in stochastic thermodynamics [28, 29].
The FEP states that, given the existence of a boundary, we can construct a map relating (the dynamics of) internal states to (those of) the environment across the boundary. In particular, if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma:\mathcal{M}\to\mathcal{N}$ such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, sending conditional modes to conditional modes [22, Lemma 4.3]. We construct this map by composing the natural projections to boundary states, $\bm{\mu}:\mathcal{B}\to\mathcal{M}$ and $\bm{\eta}:\mathcal{B}\to\mathcal{N}$, into
$\bm{\eta}\circ\bm{\mu}^{-1}:\mathcal{M}\to\mathcal{N}.$
We have assumed $\bm{\mu}$ is at minimum an injection, hence bijective on its image, from which we can take a left-cancellative pseudo-inverse. Under hypotheses on these maps, for example, that two blanket states which map to the same value of $\mu$ also map to the same value of $\eta$, this is a useful function, in the sense that it associates a unique pair of conditional modes to each blanket state. [^1] One can see [4] for a proof looking at the properties of solutions to certain linear systems of equations. Alternatively, note that this is evident from taking a tensor-Hom adjunct of the composition, $b\mapsto\xi(b,-)$, and asking that this map be an injection in the sense that when $i=j$, the map takes $\hat{\mu}_{i}=\hat{\mu}_{j}$ to equal images $\hat{\eta}_{i}=\hat{\eta}_{j}$ [22, pp 25]. Importantly, in the definition of $\sigma$, we can pass an arbitrary internal state in; its image will simply not be the conditional mode of the environment, and hence it will not parameterise a density consistent with $p(\eta\mid b)$.
We then postulate the existence of a family of probability densities parameterised by values of some known quantity, called the variational density $q(\eta;\hat{\eta})$. Generically this is constructed by taking a family of Gaussian densities with different modes. By the previous discussion, we can interpret this as the likely configuration of the environment, knowing the environment must be amenable to a certain internal state on average (given a corresponding blanket state), $q(\eta;\sigma(\mu_{b}))$. It is immediate that the probability density parameterised by the conditional mode $\hat{\eta}_{b}$ ought to be the same as the conditional probability $p(\eta\mid b)$. Moreover, if we assume the system is coupled by a particular partition, we naturally have $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$. Together, we have supposed that the object exists as the preimage of $\hat{\eta}_{b}$, or tends to do so, and hence, that the environmental states that are likely under a given conditional mode of the object ought to be equivalent to the probabilities over external states conditioned on that blanket state.
Because we are identifying a parameter value which both (i) has the greatest log-likelihood of being observed and (ii) fits a parameterised distribution over unobserved states to a distribution over unobserved states given observed states, this set up naturally has the structure of a variational inference problem [30]. As discussed, the quantity known as variational free energy scores the difference between the variational density and the true conditional density in terms of the relative entropy from one to the other, known as the KL divergence, added to the surprisal of internal and blanket states:
[…]
In the context of variational inference, we consider both terms in order to obtain a parameter for which $q=p$ almost surely and identify the least surprising (or most likely) parameter; the interpretation of the surprisal term is obvious, whilst the entropy term prevents $q$ from concentrating about the mode, spreading to match $p$ [^2] Note that as before, we can tell this story for the trajectories of a subsystem, where we take $q(\gamma(\eta_{0});u_{b})$ and $p(\gamma(\eta_{0})\mid\gamma(b_{0}))$ instead, where $u_{b}$ is the expected path given an observed trajectory of blanket states. The use of maximum entropy in the setting of sample paths is reviewed in [31]. Again, in the context of variational inference, the least surprising parameter is also the parameter with the most evidence—meaning objects that are ‘good’ at occupying their characteristic states can be described as self-evidencing. This has the interpretation of minimising the surprisal of observations of $\mu,b$ by inferring probabilities of unobserved latent states $\eta$ using the so-called discriminative model $p(\eta\mid b,\mu)$.
Importantly, the variational free energy is always greater than or equal to the surprisal, since the KL divergence is non-negative in general. Minimising the free energy effectively implements a form of inference, where we have inferred the best possible approximation to the true conditional or posterior density. That is, when the KL term is zero, the variational density and the density over external states given the state of the blanket are equal up to sets of zero probability. We observe (by properties of $\sigma$) that this is only possible when $\mu=\hat{\mu}_{b}$. In the limit of exact Bayesian inference, when the functional form of posterior beliefs encoded by internal states coincides with the posterior density over external states, the variational free energy reduces to the self-information of internal and blanket states. (Refer to e.g. [22, Lemma 4.2], or the variational free energy lemma in [27] for more discussion.)
This returns us to our earlier discussion about parameter estimation in the FEP. In this case (and again invoking the Laplace approximation), the surprisal being non-zero has a neat interpretation as the uncertainty that the object cannot resolve by being at the mode. Since the object is conditionally independent of its environment, it cannot ‘see’ beyond the Markov blanket—$N_{t}$ is an unobserved latent variable causing observed blanket states—lending this inference process a physical interpretation appropriate for self-organising and adaptive objects as well.
We will now give the following result: the object is at its most likely point in state or path space if and only if it minimises variational free energy. We will recall that the KL divergence is non-negative, implying
$D_{\mathrm{KL}}(q,p)+\mathcal{L}(\mu,b)\geq\mathcal{L}(\mu,b)$
saturating at $\mu=\hat{\mu}_{b}$ by construction. Since we defined $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$, and any movement in parameter space $\hat{\mu}_{b}+v$ adds a non-zero KL term to $\mathcal{L}(\hat{\mu}_{b},b)$, we conclude
$\nabla_{\mu}\mathcal{L}(\mu,b)=0\iff\nabla_{\mu}F(\mu,b)=0.$
What we have shown is there is a natural interpretation of the most likely state as a parameter estimation, in the sense of Bayesian inference. This argument, whilst logically straightforward, conceals a more subtle point: all of this begins from the existence of a generative model. In particular, since the joint probability $p(\eta,b,\mu)$ factorises into $p(\eta\mid b,\mu)p(b,\mu)$ (see the probabilistic chain rule), logarithms of products decompose additively, and we have $p(\eta\mid b,\mu)=p(\eta\mid b)$ by assumption, we have
[…]
In particular, if we begin with a generative model and a variational density, we are led to an inference with a satisfactory physical interpretation and explanatory power (we will dive into the latter in the section on dynamics). Not only is this the case, but by elementary mathematical considerations [^3] Notice that the support of the generative model and variational density differ. Whilst we may still take the KL divergence if the variational density is absolutely continuous with respect to the generative model, this should not be expected to be meaningful. we find we ought to take this factorisation. This observation is the main point of our proposal. The use of this form of the free energy has an even more straightforward interpretation as using an evidence lower bound to infer the generative model, such that the object can generate samples of how environmental states are likely to vary with internal states. Our objective in using the FEP is to use this mathematical fact to replace classic models of stochastic processes—which use the minimisation of surprisal—with a more tractable and interpretable quantity: the variational free energy.
2.2.1 An overview so far
All of this suggests that if a model of a subsystem occupies the states which we expect it to, given what the subsystem is, then we can model that subsystem as performing inference (see [19]). The idea is predicated on the tautology that so long as we observe that subsystem as being ‘the thing that it is’, then it must (by definition) be in a state that is amenable to its continued existence as that kind of thing. Similarly, if an object remains identifiable as a particular thing through time, it must be evolving in the way that we expect of that thing. Aaronson refers to this idea—the existence of a system in a certain state being conditioned on the result of a computation (and therefore encoding that computation in its state)—as “anthropic computation” [32]. One phrasing of our proposal is that using generative models in the study of coupled random dynamical systems is a way of leveraging anthropic computation to build models of systems in statistical physics, and that the FEP gives a formal apparatus with which we can build such models.
Put a different way, the FEP provides us with tools that allow us to model any particle, or “thing”—in the sense that it exists as a separable, identifiable thing—as if it were inferring the external states causing its boundary states, and (re)acting so as to remain in system-like configurations. This interpretation of the dynamics as inference comes from the minimisation of a free energy functional of beliefs about the causes of blanket states, as parameterised by internal states, which guarantees—or describes—the minimisation of the surprisal of $(\mu,b)$. Using this approach, we can understand the dynamics or observable behaviour of things as realising a path of minimal free energy [3, 27]. Since the object cannot ‘see’ past its Markov blanket, it must infer the external states it is interacting with; moreover, when it does so, it can be said to maximise self-evidence.
It is interesting to note that there is no guarantee the system’s model is good; in fact, we expect simple systems with poor ability to self-sustain to have simple models, meaning the FEP gives us concrete predictions about how “anthropic computers” might behave when it is too difficult to minimise a given potential. This has been discussed in the context of the FEP in [33, 34].
2.3 Dynamically
We will now explore what the FEP says about the dynamics of objects that minimise variational free energy, focusing now on the writing of a dynamical system in terms of free energy gradients—and how it relates physical properties of an object to the dynamics of beliefs parameterised by the likely states or trajectories of that object. For similar reviews one can see [2, 4, 19, 27]. We will give a quick example of how the heat dissipated into the environment by an object behind a boundary maintaining a non-equilibrium steady state can be written in terms of the variational free energy.
Assume the state space $\mathcal{N}$ is some open, flat Riemannian manifold of dimension $n$ ($\dim\mathcal{B}=d$, $\dim\mathcal{M}=m$, respectively), where we set $\ell=n+d+m$. Let $H_{t}$ be a one-dimensional standard Brownian motion: $\{H_{t}\}_{t:\tau}$ has stationary and independent increments, $H_{0}=0$ almost surely, and $t\mapsto H_{t}$ is continuous almost surely for all $t\in[0,T]$. Now let $W_{t}$ be a $k$-vector of independent Brownian motions (also called a $k$-dimensional Wiener process).
By elementary arguments in stochastic analysis, we know that we can write any stochastic differential equation (SDE) with an equilibrium measure as
[…]
Arguments in the literature imply this holds for non-equilibrium steady states as well, with the matrices described in [35, 36]:
[…]
where $\nu$ is an $\ell$-by-$k$ Lipshitz continuous matrix field with bounded linear growth. In stochastic thermodynamics the surprisal of a path given a particular ensemble of paths is often used—in our notation,
[…]
where the entries in the diffusion matrix may change in time. In each situation we have a vector field—the gradient flow of surprisal—added to a Gaussian distributed random variable with mean zero and variance $t$. (Note that from here onwards we will fix all fluctuation- and curvature-related constants to $+1$.)
Recall that $\nabla_{\mu}\mathcal{L}(\mu,b)=0$ if and only if $\nabla_{\mu}F(\mu,b)=0$. An infinitesimal variation in the KL divergence is the Fisher information; hence, using the Laplace approximation, in a neighbourhood of the mode the KL term is a constant. In that case, for a given blanket state $B_{t}=\beta$ (fixed in time) the conditional mode evolves [^4] With error bounded locally by some cubic polynomial in $(\mu-\hat{\mu}_{\beta})$. Note that since $q_{\mu}(\eta)$ is exponential by assumption, the Hessian of this free energy is non-singular and positive-definite, such that (i) second-order terms suffice to describe its behaviour in a neighbourhood of the minimum and (ii) first-order terms vanish at the minimum. as
$\partial_{t}\hat{\mu}_{\beta}(t)=-\nabla_{\mu}F(\mu,\beta)$
coming from the following SDE on the state space ××$\mathcal{M}\crossproduct\mathcal{B}\crossproduct\mathbf{R}^{2k}$:
[…]
and so on for the other equations above.
Here we recover our original notion of replacing the dynamics of the object with an inference process: the dynamics of the most likely state of the object approximately perform a minimisation of the variational free energy functional. In the path setting this becomes even more apparent, since the most likely path simply is the solution to the equations of motion minimising free energy.
2.3.1 Stochastic thermodynamics
We may now ask whether the dynamics of beliefs, seen as performing variational inference, can be related to the physical dynamics of states parameterising those beliefs. We show that we can relate performing inference to dissipating heat using this FEP.
We will consider a thermostatted system—an object embedded in a heat bath of temperature $T$ called a medium or an environment—at non-equilibrium steady state, such as a colloidal particle moving in one spatial dimension. The force on the particle is a function of both its state and a time-independent control parameter [28]. We set the blanket state as the control parameter, since this determines the mode of the non-equilibrium steady state. In this case, since we have set $\partial_{t}\hat{\mu}_{b}(t)=-\nabla F(\mu,b)$ in a neighbourhood of the mode, under mild estimates on the expectation we have the Langevin equation
$\partial_{t}\mu(t)=-\nabla_{\mu}F(\mu_{t},b)+\xi_{t}$
whose solutions for different realisations of noise are trajectories of the particle, $\gamma(\mu_{0})$. As in Freidlin–Wentzell theory, such trajectories are generated with probability
[…]
which is simply the magnitude of fluctuations (with the extra term motivated in [37]).
[…]
$\mathbf{q}[\gamma(\mu_{0})]=T\log\frac{p(\gamma(\mu_{0}))}{p(\tilde{\gamma}(\mu_{0}))}$ (1)
where $\tilde{\gamma}(\mu_{0})$ is a time-reversed path.
Computing the difference contained in (1), the heat dissipated into the environment is
[…]
This construction could be read as an implicit thermodynamics of belief updating; in the sense that it expresses the work done by a system in terms of inference. This nicely recapitulates the formulation of thermodynamics in terms of information length, based upon dynamics (of internal states) [40, 41]. See also: [4]. For example, [40] concludes that “Bennett’s classic acceptance ratio method for measuring free energy differences also measures thermodynamic length”. However, the free energy differences in this setting refer to thermodynamic free energy differences of internal states. The dénouement of the current treatment is that the same conclusions apply to variational free energy and, implicitly, beliefs about external states held by internal states.
3 Summary remarks
To summarise: formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. In the FEP, this Lyapunov function is surprisal or variational free energy, depending on the setting. In other words, to understand the claim that the FEP models a system as doing inference, it is crucial to understanding the ‘inference’ component of surprisal minimisation. Surprisal is simply the negative log probability. Variational free energy is a tractable (i.e., easily computed) upper bound on surprisal, where minimising variational free energy allows us to infer a target density by approximating it with a variational density. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised (i.e. the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics. In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [23, 42].
More specifically, suppose there exists a conditional density over external states, $q(\eta;\hat{\eta}_{b},\vartheta_{\eta\mid b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ and variance $\vartheta_{\eta\mid b}$ of the environment. For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance. This guarantees that finding the right values of $\hat{\eta}_{b},\vartheta_{\eta\mid b}$ are sufficient for $q$ to equal the posterior almost surely. More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
The FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [22, Lemma 4.3]. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$ and $[\eta-\sigma(\hat{\mu}_{b})]^{2}$, meaning that the particle models its environment. Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
If free energy (or surprisal) remains irreducibly high, it is likely that the system occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water). These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
Similarly, note the direction of the implication in Claim 1. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [22]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [19, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
From the viewpoint of model building, since internal states are by construction unobservable, we can in principle tell any story that we like about them—provided that this story is supported well enough by observable evidence. One such story is that those internal states look as if they parameterise a variational density over external states. Taking this explanation seriously, and leveraging it as an approach to modelling physical systems, is applying the FEP.
4 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. In the physics, this means that the FEP is the foundation of a method to be applied [3]. So, how might one apply the FEP?
At the most general level of analysis, it could be argued that any scientific process whatsoever is an application of the FEP: in the sense that any scientific explanation or model involves resolving uncertainty about some observed thing or dynamics, under some hypotheses about the observed world.
More specifically, how might one apply the FEP to a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: where a system that conforms to the dependence structure of the generative model is said to entail that generative model.
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. And even if one could, the internal states would be forever sequestered behind blanket states.
In short, applications of the FEP place the generative model centre stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
5 A map of that part of the territory that behaves as if it were a map
We argue that the FEP can be understood, metaphorically, a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map. In other words, the FEP provides us with tools to understand the mathematically striking property of self-organising systems: that they look as if they infer, track, reflect, or represent the statistics of their environment. Even more simply, it gives us tools to model systems that look as if they are modelling the world.
How does FEP-theoretic technology allow us to model self-organising systems as modelling their world? The generative model or joint probability $p(\eta,b,\mu)$ plays a central role. This generative model can be read as a probabilistic specification of the states or paths characteristically occupied by the joint particle-environment system. In other words, our scientific model (or map) includes the generative model entailed by the system (i.e., the territory). In this sense, the generative model we hypothesise—to account for observations of an object—allows us to represent the constraints on the kinds of states the joint system can be found in. In other words, the generative model is a modeller’s map of the whole “territory” that characterises the observed system and its environment (including the modeller, as an observer in the environment).
The technology of the FEP allows us to say that certain components of the system being considered look as if they track other components, and that this is a feature of our map contained in a generative model under a particular partition. In providing us with a calculus allowing us to model things as modelling other things, the FEP provides us with a map (a scientific model) of any possible map or tracking relationship between things (namely, a mapping between internal and external states, mediated by blanket states). [^5] Here, we are assuming representationalist accounts of scientific practice. For alternatives, see [10].
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing a Bayesian mechanical account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [43, 44]. Related to this, views of physical systems implementing computations has been articulated [45, 32], as well as views where systems self-organise by computing predictions [46, 47, 42, 29]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. See also [10].
In summary, our scientific, FEP-theoretic model is, metaphorically, a map that allows us to say that some (internal) subset of the system looks as if it possesses a map; which we interpret formally as tracking the statistics of another (external) subset. Our map (as scientists and modellers) can be identified in an unproblematic way as the generative model of the system, which FEP-theoretic technology enables us to write down. It is by construction that our model is a model of the modelling capacities of subsets of the real-world system: that is, our map is precisely a map of how objects behave as if they were maps, allowing us to construct scientific models of the maps encoded (or looking as if they were encoded) by the internal subset of the system considered. This is an answer to allegations that FEP-theoretic modelling commits the map-territory fallacy: namely, it shows that there is no conflation of the map and the territory, in the sense of the map we postulate as being embodied by the system within the mathematical machinery of the FEP.
Taking all the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system. This echoes seminal work in philosophy by Wittgenstein [48]. In his famous Tractatus Logico-Philosophicus, Wittgenstein set out to delimit the domain of sensible propositions by determining the general form of a proposition, an utterance able to carry a truth value. This set ultimate constraints on what is sensible, and bounded what is meaningful “from inside” of language itself. Similarly, the FEP provides ultimate constraints on possible maps or modelling relations, which arise from what it means to be a map or model of a physical process at all. Indeed, any modelling process consistent with the laws of physics must conform to the FEP, such that the FEP sets ultimate limits on what can count as a map or model—starting from within the technology of mapping or modelling itself.
In sum, the generative model is a kind of scientific model, which harnesses what we know about the dependency structure of the system. A variational density, on the other hand, is a probability density over external states or paths, implemented in terms of the sufficient statistics of external states given blanket states. It is an explicitly inferential representation of the manner in which subsets of a system with a particular partition track each other: in which case, we are describing a part of the territory that, as seen in our map, also behaves as if it were a map.
6 Some remarks on terminological idiosyncrasies of the FEP literature
In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [8, 9]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled. This allegation constitutes a fallacy, which one might call the map-territory fallacy fallacy.
Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling. Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not. In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [13]. The It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature. In fairness, these ambiguities and idiosyncrasies may lend themselves to misinterpretation. We now discuss two such cases.
6.1 Generative models and variational densities versus recognition densities
The first concerns the ambiguous discussion or description of two core constructs in the FEP literature—namely, the generative model and the variational (or recognition) density. It is noteworthy that the manner in which these constructs are deployed in the FEP-theoretic formulation is markedly different from previous Bayesian approaches. This has seldom been noted in the literature (with a few exceptions, e.g., [13]); and even when it is discussed, additional issues tend to undermine the clarity of the exposition (as noted by [9]).
In more traditional machine learning approaches, e.g., in the literature on the Helmholtz machine [49], the generative model and the recognition model correspond to the feedforward and the feedback passes that together constitute the process of data-driven Bayesian inference. In these settings, both the generative model and the recognition model are straightforwardly encoded in the connection weights of the Helmholtz machine: the generative model is encoded in the weighting of the backwards connections, which establish an information channel from high-level latent spaces of the neural network to its sensory layers; whereas the recognition model is encoded in the feedforward connections. In this setting, one assumes that any system that implements this pair of functionally differentiated weightings (e.g., a brain) implements a generative and recognition model pair.
Much of the literature on the FEP is based on this reading of the generative and recognition models: it is often said within this literature that FEP-theoretic generative models are encoded by the brain, much like these would be in a Helmholtz machine. However, whilst generative and variational models are also used in the FEP-theoretic formulation, they are deployed in a markedly different way, which we discussed extensively above.
Overlooking these differences may have contributed to the impression that theorists of the FEP conflate the metaphorical map and the territory. Recall that a generative model, generically, is just a joint probability density defined over the states (or paths) of a system. In other words, it is our map or representation of the system being modelled and its mechanics; and more particularly, it is our model of the dependencies between the variables of the system, providing us with equations of motion for the whole system (i.e., it is defined over internal, blanket, and external states/paths).
Technically, we say that the system that we are modelling entails a generative model; whereas the internal states of the particle (look, to the modeller, as if they) estimate the parameters of a conditional density over external states. This difference is crucial to understanding what is at stake in FEP-theoretic modelling: we say that one subset of variables over which the generative model is defined (i.e., the internal states) encodes a variational or recognition model over another subset (the external states). Entailment is the property that, metaphorically, “the part of the territory that behaves like a map” (the internal states that encode the parameters of the variational density) is itself part of the territory described by our map (the generative model) and so is subject to the mechanics of that territory, as described by our map. Entailment means that the relationships that hold between variables, as we represent in the generative model, are obtained as a description of the dynamics of the real-world system.
The subtle but key point is that the generative model is not encoded by the internal states of the particle being considered: it is not some kind of internal model. Rather, the generative model is a description of the dependence relations that obtain in the dynamics of a system: i.e., such-and-such dynamics entail such-and-such equations of motion and dependencies. FEP-theoretic modelling does not assume that the generative model is encoded in the brain. Arguably, to claim this would indeed constitute a map-territory fallacy. Under the FEP, parts of the system itself (i.e., by internal states) look to an external observer as if, or can be modelled as if, they encode the parameters of a variational model.
However, this raises a question: if it is our scientific model, how can the generative model play a role in inference if only a conditional posterior model is encoded by the system? To claim that a system uses or leverages its generative model is, in some sense, a shorthand expressing the fact that the system leverages its own causal structure, and the data patterns that it elicits through action, to generate patterns of adaptive behaviour—all of which are subject to the regularities that are captured in the generative model. This is a statement of embodied and enactive cognition [50, 13]. The system, in having sensory states, also has access to the causal structure in the generative model.
So, the generative model is our model or map, but it denotes the causal structure of the system being modelled (i.e., the territory). This causal structure—and in particular, the presence of a Markov blanket or particular partition—means that some parts of the system (the internal states) encode a variational density over some other parts of the system (the external states); that is, they “act like a map” or model the statistics of external states.
6.2 Generative models and generative processes
The second ambiguity concerns the generative model and the so-called “generative process”. Some presentations of the FEP sometimes define the generative model as a scientific model of some data-generating generative process; see, for instance, [51, 13]). On this reading, the generative model is our “map” of some real-world process, namely, the generative process—which would then correspond to the metaphorical territory.
In practice, however, in FEP-theoretic modelling, this distinction does not map onto that between a scientific model and real-world system, i.e., onto the distinction between the map and territory. Perhaps confusingly, the generative process does not denote (the structure of) the real data-generating process that we are trying to model. In FEP-theoretic formulations, the generative process is also part of our scientific model.
It should be noted that, practically, the distinction between generative model and generative process only comes into play when we are implementing the dynamics of a specific system in simulations. When we write code to simulate the system of interest, typically, we need to specify (i) the dependencies that hold between the variables of a system, which we call the generative model; and (ii) the equations of motion that govern the time evolution of the external states per se in our simulation. The latter is labelled the ‘generative process’, since it generates the dynamics of external states, which are tracked by internal states. The key thing to note, however, is that generative process is also part of our scientific model—it harnesses what we know about the physics of the situation described by the generative model, in which the thing is embedded and ‘lives’.
7 Conclusion
We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle. We have further argued that this allegation itself constitutes a map-territory fallacy fallacy. In distinguishing the generative model and the variational density—and the generative model and the generative process from the real-world system that we aim to model—the technology of FEP-theoretic modelling allows us to construct a map of that part of the territory which looks to an external observer as if or behaves as if it were a map, without committing model reification. The FEP is thus, metaphorically, a map of any possible map whatsoever of, or held by, a physical system. One ought to celebrate this ‘territory-map mapping,’ and the FEP is ultimately, at its core, a principled approach to the formalisation of this mapping.

### v2

@@
-An approach to non-equilibrium statistical physics using variational Bayesian inference
+A framework for the use of generative modelling in non-equilibrium statistical mechanics
 Abstract
@@
 This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic.
-The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna and the Laufer Centre for Physical and Quantitative Biology, and the informal discussions that followed.
+The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed.
+We thank the organisers and audiences of both.
+DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York.
 1 Introduction
@@
 In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt.
-One approach to making such models, using variational (approximate Bayesian) inference [2, 3], describes how a physical object reflects the properties of its environment and vice versa.
-Bayesian mechanics under the free energy principle (FEP) [4, 5] aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference), resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled.
+One approach to making such models, using variational (approximate Bayesian) inference [2, 3, 4], describes how a physical object reflects the properties of its environment and vice versa.
 In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena.
-More specifically, we argue that generative modelling—premised on the FEP—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action.
-We use the FEP to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
-We identify a particular issue, which the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way.
-Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that the FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled).
+More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [5]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action.
+We use the ensuing variational free energy principle to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
+Our proposed free energy principle (FEP) has been explored in some depth in prior literature, from the viewpoint of modelling control systems using the property that they generally look as if they track or infer the statistics of perturbations to their setpoints.
+In this paper we will make a more general case for its utility in statistical mechanics at large.
+After reviewing prior work on the FEP, we identify a particular issue that the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way, offering advantages in terms of tractability and explanation, by using gradients of variational free energy to describe the dynamics of interacting systems.
+Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled).
 The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [6, 7].
-Some have referenced this fallacy in relation to the FEP [8, 9].
+Some have referenced this fallacy in relation to prior work on the FEP [8, 9].
 In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled.
-(See [10, 11, 12, 3] for related critical discussions of such claims.)
-According to this line of thinking, using Bayesian mechanics to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification.
-The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific FEP-theoretic or Bayesian mechanical model) with the “territory” (the real-world system itself).
+(See [10, 11, 12, 3] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification.
+The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific, FEP-theoretic model) with the “territory” (the real-world system itself).
 This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [10, 11].
-By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of anything whatsoever that exists in the physical world.
-Here, we discuss the use of the FEP in statistics and statistical physics, and what advantages it offers; we then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished.
+By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
+After describing the advantages of this FEP we will then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished.
 We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy.
@@
 We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object.
-That is to say, a model of an object can always be written as modelling or inferring the structure of its environment.
+That is to say, a model of an object can always be written as if it were modelling or inferring the structure of its environment.
 To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
@@
 The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket.
-Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is (a form of) Bayesian inference.
+Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
 We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
@@
 Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
-We first present a conceptual, and then technical, introduction to the FEP and Bayesian mechanics—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients.
+The structure of the paper is as follows.
+We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients.
 This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer.
@@
 We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other.
-We propose that the FEP and Bayesian mechanics provide a map of any possible map whatsoever—of, or held by—a physical system.
-(This includes the use of the FEP itself by a modeller.) We suggest that this leads, in a Wittgensteinian manner, to constraints on all possible maps, or models of physical things.
-This is meant in terms of logically necessary preconditions for sense-making, and thus the parsing of sensory streams on which cognition is contingent [15].
-In particular, these constraints arise from what it means to be a map or model of a physical process at all.
+In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system.
+(This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all.
+In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [20, 21], the FEP is a natural expression of statistical physics as a process of making models.
 We remark that we will not claim to invalidate the map-territory fallacy itself.
 That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]).
-Models built using the FEP are still scientific models, and the properties of FEP theoretic model also run the risk of reification.
+Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification.
 Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
-2 The free energy principle
-2.1 Conceptually
-In this subsection we will provide an up-to-date overview of the concepts behind the FEP.
-One can also see [3] for a discussion of the main ideas.
-See the following subsection for more mathematical details.
-Conceptually, the FEP starts from a partitions an overall “system” into three component sets of states: the internal states $\mu$ of a ‘thing’ or ‘particle’, an environment (which possesses external states $\eta$, relative to the thing in question), and an interface separating but coupling the two (the so-called Markov blanket, $b$), which mediates the influence of $\eta$ on $\mu$ and vice-versa.
-The states of the system will be tuples $x\coloneqq(\eta,b,\mu)$, where $\eta$ and $\mu$ are states of two subsystems and $b$ is some locus of states such that $p(\eta\mid\mu,b)=p(\eta\mid b)$ and vice versa.
-“Particular states” are the internal states of each thing and their respective Markov blanket.
-Of note, the particular partition is akin to choosing a set of coarse-grained variables at a superordinate scale, associating each state or path at the subordinate scale to a particle at the superordinate scale (as either an internal or a blanket state), in terms of the relations of conditional independence that obtain between states or paths [20, 2].
-More precisely, we will consider systems of interacting objects, such as an open random dynamical system coupled to its environment.
-Fixing a final time $\tau$, the joint system will be denoted by the time-indexed family of random variables $\{X\}_{t:\tau}$ taking values $x\in\mathcal{X}$ at each $t$, with a specific realisation of the system (meaning, a sample path or trajectory beginning at some fixed $X_{0}=x_{0}$) being the sequence $\gamma(x_{0})=\{x_{0},x_{1},\ldots,x_{\tau}\}$.
-We will define $\gamma(x_{0})\mapsto\gamma_{t}(x_{0})$ as the projection to the state at time $t$.
-Then $B_{t}$ is what we have called the Markov blanket in $\{X\}_{t:\tau}$ at time $t$, separating it into statistically distinct subsystems (which one can then argue do not mix into each other [18]).
-We will repeat this notation for
-$\gamma_{t}(\eta_{0})=N_{t}=\eta\in\mathcal{N}$
-$\gamma_{t}(b_{0})=B_{t}=b\in\mathcal{B}$
-$\gamma_{t}(\mu_{0})=M_{t}=\mu\in\mathcal{M}$
-as well as for conditional trajectories such as
-$\gamma_{t}(\eta_{0}\mid B_{t}=b)=\eta_{b}\in\mathcal{N}\crossproduct\{b\}.$
-Although we will not go into the theory of random dynamical systems, when equipped with a time evolution map acting on the sample space, these are all the ingredients necessary to define a random dynamical system [21], and this is implicit in our discussion; for details see the account in [14].
-When $\{M\}_{t:\tau}$ is an open system coupled to an environment with states $\eta$, a value $\mu$ is called an internal state, and the subsystem $\{M\}_{t:\tau}$ is variously called an object, a thing, or a particle in the literature.
-This is the convention we will stick to.
-We can guarantee that the dynamics of a system can be described as a flow on surprisal, and thereby, cast as implementing a form of inference, whenever there exists a “particular partition” of that system.
-At the core of the FEP is a mathematical statement about the behaviour of such systems.
-For a comprehensive overview, see [3].
-In particular, the FEP allows us relate the dynamics of states or paths that are internal to the object, $\mu$, to the dynamics of external states, conditioned on boundary or blanket states $p(\eta\mid b)$—that is, understanding what goes on inside a subsystem by understanding what it is coupled to, as ‘seen’ via blanket states [5, 22].
-The logic of the FEP can be summarised as follows.
-If there exists a particular partition of a system, then there exists a coupling between internal and external states or paths, given blanket states or paths; then, since the internal dynamics of a given object are coupled to other (external) subsets of the system—in such a way that they can be read as parameters for probability densities over the external subset—the FEP describes these interacting systems using variational inference.
-This is a core aspect of the FEP that distinguishes it from a straightforward application of density dynamics to the internal states per se.
-This is where the interpretation of dynamics as a form of inference comes into play. Indeed, one way to state the FEP is:
-Claim 1.
-If a random dynamical system (i) is subject to a specific kind of sparse coupling or boundary condition, such that it possesses a Markov blanket, and (ii) has a subsystem $\{M\}_{t:\tau}$ under that blanket with characteristic states (in the sense that it occupies a pullback attractor), then $\{M\}_{t:\tau}$ can always be described as if it minimises a variational free energy functional of densities over states or trajectories of $\{N\}_{t:\tau}$ [14, 22].
-Another statement one can make is that Bayesian mechanics formalises the truism that, on average, an object will do whatever it is that it does on average, given the sort of object it is.
-Behind this tautological statement, we find a formal apparatus that enables us to model a variety of dynamical systems in a parsimonious manner.
-We hope this will allow us to take some first steps towards a mechanistic description of how self-organisation manifests in physical systems with such pullback attractors [3].
-Bayesian mechanics explains why every physical thing looks as if it were performing inference over the causes of its perturbations, forming (Bayes) optimal beliefs about those causes [22].
-Bayesian mechanics begins from the minimisation of self-information (also called surprisal) or free energy, and allows us to write down mechanical theories (i.e., laws of motion) yielding the dynamics of systems on an attractor [19].
-Surprisal is simply the negative log-probability of an observed outcome.
-To say that an object minimises (or follows a path of least) surprisal is to say that an object evolves towards the outcomes which are most probable, given the sort of thing that it is (meaning, its characteristic states of being, or characteristic trajectories through its state space).
-Interestingly, this means that surprisal is suited to act as a kind of ‘ontological potential’ for the object, in the sense that the surprisal is equivalent to a constraint on what states and outcomes are accessible to the object; the trajectories of a physical system are random curves through the gradient of surprisal over the state space, which satisfy an equation of motion derived from the surprisal itself [4, 23].
-As has been asserted in recent literature [19], one reading of the FEP would hold that the crux of the ubiquity of the FEP is that everything which is coupled to another thing holds a model of that thing to which it is coupled in the following, trivial sense: in virtue of this coupling, one subsystem estimates the parameters of the probability density over states (or trajectories) of the other subsystems (and vice versa).
-Using this property, they can be written as performing inference in the sense of variational Bayesian inference.
-This coupling is sometimes discussed in terms of the ‘probabilistic beliefs’ that one object holds about its environment.
-Given that the term ‘belief’ is quite theoretically loaded (especially in philosophy), we ought perhaps to say ‘estimator,’ instead.
-Once more, our contention is that anything which estimates the parameters of a density is engaged in a form of Bayesian inference, by mathematical definition; i.e., it is mathematically indisputable that systems that estimate statistics represent, or encode, or otherwise hold probabilistic beliefs—whether the practitioner likes the word ‘beliefs’ or not, and irrespective of what commitments one has with respect to notions such as encoding and representation (e.g., [24]).
-Generative models and variational methods allow us to licence these claims [25, 26].
-Generative models are joint probability densities over causes and consequences which, here, are used to capture the dependencies between the variables constituting $\{X\}_{t:\tau}$, meaning the way that changes in some variables cause changes in others.
-Writing down a generative model for some real object in the world (the metaphorical territory) is building a metaphorical map of that territory (a formalisation or codification of the relations between the things in that territory).
-Variational techniques comprise a family of inference techniques which allow a modeller to approximate some (usually intractable) posterior probability density under a (usually specified) generative model.
-In variational inference, we write down a parameterised probability density function called a variational density, which is used as an approximation to the posterior probability density.
-The quantity called variational free energy scores the difference between the variational density and the posterior over external states—here, a conditional density, given a particular blanket state.
-Hence, when some notion of the difference between them is zero, variational free energy can be used as a proxy for surprise, and we can approximate the posterior density; which is a form of parameterised Bayesian inference.
-Note that any given object cannot change the generative model associated to that object.
-This follows because the object can only change the variational density encoded by its internal states.
-However, to explain these changes—under a variational principle of least action—one needs to specify the generative model under which the posterior is approximated.
-This means the generative model explains the dynamics of something that is observed in a particular way.
-This is usually expressed by saying that the generative model is entailed by the thing in question.
-That an object entails a generative model is consistent with statements like “an object is a model of its environment”: e.g., in the sense of the good regulator theorem [19]).
-This follows because the generative model just is a probabilistic description of the joint states an object and its environment occupy.
-2.2 Mathematically
-With that conceptual overview being stated, we will now provide some further mathematical details.
-Similar (contemporary) reviews can be found in [4, 5, 19, 27].
-We will argue that coupled random dynamical systems partitioned under a Markov blanket naturally have the form of variational Bayesian inference problems, in the sense that we obtain the values of parameters from the minimisation of free energy or surprisal.
-This is what licences the claim that minimising free energy implements a form of inference, in which case the object models its environment.
-Ultimately this simply means that, in virtue of being coupled via blanket states to its environment, the object reflects the structure of, and data about (i.e., models), its environment—or, what it ‘believes’ that its environment is like (in the sense of being a statistical estimator, as defined above).
-Appealing to the Laplace approximation (that if a density has a unique maximum, in a neighbourhood of the most likely state, then the density is approximately Gaussian with that mean), we will suppose all random variables and all trajectories are jointly Gaussian.
-We will call the conditionally expected state—that is, the expected state given some fixed blanket state—the conditional mode $\hat{\mu}_{b}$ or $\hat{\eta}_{b}$.
-Let $\mathcal{L}(\mu,b)=-\log p(\mu,b)$ be the surprisal of internal and blanket states, such that $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$ and thus satisfies $\nabla_{\mu}\mathcal{L}(\mu,b)=0$.
-By the Freidlin–Wentzell theorem, we can argue the same is true of paths (see [19]).
-In that setting, the interpretation of surprisal as the fluctuations of a random dynamical system away from its expected path plays a crucial role in stochastic thermodynamics [28, 29].
-The FEP states that, given the existence of a boundary, we can construct a map relating (the dynamics of) internal states to (those of) the environment across the boundary.
-In particular, if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma:\mathcal{M}\to\mathcal{N}$ such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, sending conditional modes to conditional modes [22, Lemma 4.3].
-We construct this map by composing the natural projections to boundary states, $\bm{\mu}:\mathcal{B}\to\mathcal{M}$ and $\bm{\eta}:\mathcal{B}\to\mathcal{N}$, into
-$\bm{\eta}\circ\bm{\mu}^{-1}:\mathcal{M}\to\mathcal{N}.$
-We have assumed $\bm{\mu}$ is at minimum an injection, hence bijective on its image, from which we can take a left-cancellative pseudo-inverse.
-Under hypotheses on these maps, for example, that two blanket states which map to the same value of $\mu$ also map to the same value of $\eta$, this is a useful function, in the sense that it associates a unique pair of conditional modes to each blanket state.
-[^1] One can see [4] for a proof looking at the properties of solutions to certain linear systems of equations.
-Alternatively, note that this is evident from taking a tensor-Hom adjunct of the composition, $b\mapsto\xi(b,-)$, and asking that this map be an injection in the sense that when $i=j$, the map takes $\hat{\mu}_{i}=\hat{\mu}_{j}$ to equal images $\hat{\eta}_{i}=\hat{\eta}_{j}$ [22, pp 25].
-Importantly, in the definition of $\sigma$, we can pass an arbitrary internal state in; its image will simply not be the conditional mode of the environment, and hence it will not parameterise a density consistent with $p(\eta\mid b)$.
-We then postulate the existence of a family of probability densities parameterised by values of some known quantity, called the variational density $q(\eta;\hat{\eta})$.
-Generically this is constructed by taking a family of Gaussian densities with different modes.
-By the previous discussion, we can interpret this as the likely configuration of the environment, knowing the environment must be amenable to a certain internal state on average (given a corresponding blanket state), $q(\eta;\sigma(\mu_{b}))$.
-It is immediate that the probability density parameterised by the conditional mode $\hat{\eta}_{b}$ ought to be the same as the conditional probability $p(\eta\mid b)$.
-Moreover, if we assume the system is coupled by a particular partition, we naturally have $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$.
-Together, we have supposed that the object exists as the preimage of $\hat{\eta}_{b}$, or tends to do so, and hence, that the environmental states that are likely under a given conditional mode of the object ought to be equivalent to the probabilities over external states conditioned on that blanket state.
-Because we are identifying a parameter value which both (i) has the greatest log-likelihood of being observed and (ii) fits a parameterised distribution over unobserved states to a distribution over unobserved states given observed states, this set up naturally has the structure of a variational inference problem [30].
-As discussed, the quantity known as variational free energy scores the difference between the variational density and the true conditional density in terms of the relative entropy from one to the other, known as the KL divergence, added to the surprisal of internal and blanket states:
-[…]
-In the context of variational inference, we consider both terms in order to obtain a parameter for which $q=p$ almost surely and identify the least surprising (or most likely) parameter; the interpretation of the surprisal term is obvious, whilst the entropy term prevents $q$ from concentrating about the mode, spreading to match $p$ [^2] Note that as before, we can tell this story for the trajectories of a subsystem, where we take $q(\gamma(\eta_{0});u_{b})$ and $p(\gamma(\eta_{0})\mid\gamma(b_{0}))$ instead, where $u_{b}$ is the expected path given an observed trajectory of blanket states.
-The use of maximum entropy in the setting of sample paths is reviewed in [31].
-Again, in the context of variational inference, the least surprising parameter is also the parameter with the most evidence—meaning objects that are ‘good’ at occupying their characteristic states can be described as self-evidencing.
-This has the interpretation of minimising the surprisal of observations of $\mu,b$ by inferring probabilities of unobserved latent states $\eta$ using the so-called discriminative model $p(\eta\mid b,\mu)$.
-Importantly, the variational free energy is always greater than or equal to the surprisal, since the KL divergence is non-negative in general.
-Minimising the free energy effectively implements a form of inference, where we have inferred the best possible approximation to the true conditional or posterior density.
-That is, when the KL term is zero, the variational density and the density over external states given the state of the blanket are equal up to sets of zero probability.
-We observe (by properties of $\sigma$) that this is only possible when $\mu=\hat{\mu}_{b}$.
-In the limit of exact Bayesian inference, when the functional form of posterior beliefs encoded by internal states coincides with the posterior density over external states, the variational free energy reduces to the self-information of internal and blanket states.
-(Refer to e.g.
-[22, Lemma 4.2], or the variational free energy lemma in [27] for more discussion.)
-This returns us to our earlier discussion about parameter estimation in the FEP.
-In this case (and again invoking the Laplace approximation), the surprisal being non-zero has a neat interpretation as the uncertainty that the object cannot resolve by being at the mode.
-Since the object is conditionally independent of its environment, it cannot ‘see’ beyond the Markov blanket—$N_{t}$ is an unobserved latent variable causing observed blanket states—lending this inference process a physical interpretation appropriate for self-organising and adaptive objects as well.
-We will now give the following result: the object is at its most likely point in state or path space if and only if it minimises variational free energy.
-We will recall that the KL divergence is non-negative, implying
-$D_{\mathrm{KL}}(q,p)+\mathcal{L}(\mu,b)\geq\mathcal{L}(\mu,b)$
-saturating at $\mu=\hat{\mu}_{b}$ by construction.
-Since we defined $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$, and any movement in parameter space $\hat{\mu}_{b}+v$ adds a non-zero KL term to $\mathcal{L}(\hat{\mu}_{b},b)$, we conclude
-$\nabla_{\mu}\mathcal{L}(\mu,b)=0\iff\nabla_{\mu}F(\mu,b)=0.$
-What we have shown is there is a natural interpretation of the most likely state as a parameter estimation, in the sense of Bayesian inference.
-This argument, whilst logically straightforward, conceals a more subtle point: all of this begins from the existence of a generative model.
-In particular, since the joint probability $p(\eta,b,\mu)$ factorises into $p(\eta\mid b,\mu)p(b,\mu)$ (see the probabilistic chain rule), logarithms of products decompose additively, and we have $p(\eta\mid b,\mu)=p(\eta\mid b)$ by assumption, we have
-[…]
-[…]
-In particular, if we begin with a generative model and a variational density, we are led to an inference with a satisfactory physical interpretation and explanatory power (we will dive into the latter in the section on dynamics).
-Not only is this the case, but by elementary mathematical considerations [^3] Notice that the support of the generative model and variational density differ.
-Whilst we may still take the KL divergence if the variational density is absolutely continuous with respect to the generative model, this should not be expected to be meaningful.
-we find we ought to take this factorisation.
-This observation is the main point of our proposal.
-The use of this form of the free energy has an even more straightforward interpretation as using an evidence lower bound to infer the generative model, such that the object can generate samples of how environmental states are likely to vary with internal states.
-Our objective in using the FEP is to use this mathematical fact to replace classic models of stochastic processes—which use the minimisation of surprisal—with a more tractable and interpretable quantity: the variational free energy.
-2.2.1 An overview so far
-All of this suggests that if a model of a subsystem occupies the states which we expect it to, given what the subsystem is, then we can model that subsystem as performing inference (see [19]).
-The idea is predicated on the tautology that so long as we observe that subsystem as being ‘the thing that it is’, then it must (by definition) be in a state that is amenable to its continued existence as that kind of thing.
-Similarly, if an object remains identifiable as a particular thing through time, it must be evolving in the way that we expect of that thing.
-Aaronson refers to this idea—the existence of a system in a certain state being conditioned on the result of a computation (and therefore encoding that computation in its state)—as “anthropic computation” [32].
-One phrasing of our proposal is that using generative models in the study of coupled random dynamical systems is a way of leveraging anthropic computation to build models of systems in statistical physics, and that the FEP gives a formal apparatus with which we can build such models.
-Put a different way, the FEP provides us with tools that allow us to model any particle, or “thing”—in the sense that it exists as a separable, identifiable thing—as if it were inferring the external states causing its boundary states, and (re)acting so as to remain in system-like configurations.
-This interpretation of the dynamics as inference comes from the minimisation of a free energy functional of beliefs about the causes of blanket states, as parameterised by internal states, which guarantees—or describes—the minimisation of the surprisal of $(\mu,b)$.
-Using this approach, we can understand the dynamics or observable behaviour of things as realising a path of minimal free energy [3, 27].
-Since the object cannot ‘see’ past its Markov blanket, it must infer the external states it is interacting with; moreover, when it does so, it can be said to maximise self-evidence.
-It is interesting to note that there is no guarantee the system’s model is good; in fact, we expect simple systems with poor ability to self-sustain to have simple models, meaning the FEP gives us concrete predictions about how “anthropic computers” might behave when it is too difficult to minimise a given potential.
-This has been discussed in the context of the FEP in [33, 34].
-2.3 Dynamically
-We will now explore what the FEP says about the dynamics of objects that minimise variational free energy, focusing now on the writing of a dynamical system in terms of free energy gradients—and how it relates physical properties of an object to the dynamics of beliefs parameterised by the likely states or trajectories of that object.
-For similar reviews one can see [2, 4, 19, 27].
-We will give a quick example of how the heat dissipated into the environment by an object behind a boundary maintaining a non-equilibrium steady state can be written in terms of the variational free energy.
-Assume the state space $\mathcal{N}$ is some open, flat Riemannian manifold of dimension $n$ ($\dim\mathcal{B}=d$, $\dim\mathcal{M}=m$, respectively), where we set $\ell=n+d+m$.
-Let $H_{t}$ be a one-dimensional standard Brownian motion: $\{H_{t}\}_{t:\tau}$ has stationary and independent increments, $H_{0}=0$ almost surely, and $t\mapsto H_{t}$ is continuous almost surely for all $t\in[0,T]$.
-Now let $W_{t}$ be a $k$-vector of independent Brownian motions (also called a $k$-dimensional Wiener process).
-By elementary arguments in stochastic analysis, we know that we can write any stochastic differential equation (SDE) with an equilibrium measure as
-[…]
-Arguments in the literature imply this holds for non-equilibrium steady states as well, with the matrices described in [35, 36]:
-[…]
-where $\nu$ is an $\ell$-by-$k$ Lipshitz continuous matrix field with bounded linear growth.
-In stochastic thermodynamics the surprisal of a path given a particular ensemble of paths is often used—in our notation,
-[…]
-where the entries in the diffusion matrix may change in time.
-In each situation we have a vector field—the gradient flow of surprisal—added to a Gaussian distributed random variable with mean zero and variance $t$.
-(Note that from here onwards we will fix all fluctuation- and curvature-related constants to $+1$.)
-Recall that $\nabla_{\mu}\mathcal{L}(\mu,b)=0$ if and only if $\nabla_{\mu}F(\mu,b)=0$.
-An infinitesimal variation in the KL divergence is the Fisher information; hence, using the Laplace approximation, in a neighbourhood of the mode the KL term is a constant.
-In that case, for a given blanket state $B_{t}=\beta$ (fixed in time) the conditional mode evolves [^4] With error bounded locally by some cubic polynomial in $(\mu-\hat{\mu}_{\beta})$.
-Note that since $q_{\mu}(\eta)$ is exponential by assumption, the Hessian of this free energy is non-singular and positive-definite, such that (i) second-order terms suffice to describe its behaviour in a neighbourhood of the minimum and (ii) first-order terms vanish at the minimum.
-as
-$\partial_{t}\hat{\mu}_{\beta}(t)=-\nabla_{\mu}F(\mu,\beta)$
-coming from the following SDE on the state space ××$\mathcal{M}\crossproduct\mathcal{B}\crossproduct\mathbf{R}^{2k}$:
-[…]
-and so on for the other equations above.
-Here we recover our original notion of replacing the dynamics of the object with an inference process: the dynamics of the most likely state of the object approximately perform a minimisation of the variational free energy functional.
-In the path setting this becomes even more apparent, since the most likely path simply is the solution to the equations of motion minimising free energy.
-2.3.1 Stochastic thermodynamics
-We may now ask whether the dynamics of beliefs, seen as performing variational inference, can be related to the physical dynamics of states parameterising those beliefs.
-We show that we can relate performing inference to dissipating heat using this FEP.
-We will consider a thermostatted system—an object embedded in a heat bath of temperature $T$ called a medium or an environment—at non-equilibrium steady state, such as a colloidal particle moving in one spatial dimension.
-The force on the particle is a function of both its state and a time-independent control parameter [28].
-We set the blanket state as the control parameter, since this determines the mode of the non-equilibrium steady state.
-In this case, since we have set $\partial_{t}\hat{\mu}_{b}(t)=-\nabla F(\mu,b)$ in a neighbourhood of the mode, under mild estimates on the expectation we have the Langevin equation
-$\partial_{t}\mu(t)=-\nabla_{\mu}F(\mu_{t},b)+\xi_{t}$
-whose solutions for different realisations of noise are trajectories of the particle, $\gamma(\mu_{0})$. As in Freidlin–Wentzell theory, such trajectories are generated with probability
-[…]
-which is simply the magnitude of fluctuations (with the extra term motivated in [37]).
-[…]
-$\mathbf{q}[\gamma(\mu_{0})]=T\log\frac{p(\gamma(\mu_{0}))}{p(\tilde{\gamma}(\mu_{0}))}$ (1)
-where $\tilde{\gamma}(\mu_{0})$ is a time-reversed path.
-Computing the difference contained in (1), the heat dissipated into the environment is
-[…]
-[…]
-This construction could be read as an implicit thermodynamics of belief updating; in the sense that it expresses the work done by a system in terms of inference.
-This nicely recapitulates the formulation of thermodynamics in terms of information length, based upon dynamics (of internal states) [40, 41].
-See also: [4].
-For example, [40] concludes that “Bennett’s classic acceptance ratio method for measuring free energy differences also measures thermodynamic length”.
-However, the free energy differences in this setting refer to thermodynamic free energy differences of internal states.
-The dénouement of the current treatment is that the same conclusions apply to variational free energy and, implicitly, beliefs about external states held by internal states.
-3 Summary remarks
-To summarise: formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system.
+2 Summary remarks on the free energy principle
+Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [19, 3, 4] for a more extended look).
+The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [22], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 23, 19].
+Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system.
 The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function.
@@
 The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration.
-Minimising surprisal implies that the variational free energy is minimised (i.e.
-the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
-The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics.
+Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
+We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state distribution $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$.
+After Pearl [17] we refer to $B$ as the Markov blanket of the system.
+We then postulate a conditional density over external states, $q(\eta;\hat{\eta}_{b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ of the environment.
+For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance.
+This guarantees that finding the right value of $\hat{\eta}_{b}$ is sufficient for $q$ to equal the posterior almost surely.
+More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
+The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g.
+Freidlin–Wentzell.
 In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal.
 This gradient flow is what we identify as actually being the process (i.e., the performance) of inference.
-Explicit examples of this can be found in [23, 42].
-More specifically, suppose there exists a conditional density over external states, $q(\eta;\hat{\eta}_{b},\vartheta_{\eta\mid b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ and variance $\vartheta_{\eta\mid b}$ of the environment.
-For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance.
-This guarantees that finding the right values of $\hat{\eta}_{b},\vartheta_{\eta\mid b}$ are sufficient for $q$ to equal the posterior almost surely.
-More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
-The FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [22, Lemma 4.3].
-In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$ and $[\eta-\sigma(\hat{\mu}_{b})]^{2}$, meaning that the particle models its environment.
+Explicit examples of this can be found in [24, 25, 19].
+More specifically, the FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [23, Lemma 4.3].
+In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution.
 Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like.
@@
 If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
-If free energy (or surprisal) remains irreducibly high, it is likely that the system occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water).
+Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy
+[…]
+Of physical interest [^1] See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics.
+is the observation that the divergence term is non-negative, meaning $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
+$F(\mu,b)=D_{\mathrm{KL}}(q(\eta;\sigma(\mu))\|p(\eta\mid b))-\log p(\mu,b)\geqslant-\log p(\mu,b)$
+with equality precisely when the KL term vanishes. From that reasoning, writing
+[…]
+in a neighbourhood of the mode we can equivalently write
+[…]
+Though seemingly vacuous, this move opens a world of interesting interpretations for models built according to this reasoning.
+For example: by existing a particular way inside the joint state space, the particle can be seen as encoding (via its mode) an inference about a generative model over the entire joint state space, resulting from carving out some ‘niche’ within it.
+That is to say, the environment must likely be some certain way for the system to exist the way it does.
+On the other hand, if free energy (or surprisal) remains irreducibly high, it is likely that the particle occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water).
+Correspondingly, its inference about the generative model would be incorrect, concentrating mass in the wrong areas of the joint state space.
 These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
-Similarly, note the direction of the implication in Claim 1.
+Similarly, note the direction of the implication in our claims.
 In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [23].
@@
 In the physics, this means that the FEP is the foundation of a method to be applied [3].
-So, how might one apply the FEP?
-At the most general level of analysis, it could be argued that any scientific process whatsoever is an application of the FEP: in the sense that any scientific explanation or model involves resolving uncertainty about some observed thing or dynamics, under some hypotheses about the observed world.
-More specifically, how might one apply the FEP to a particular particle, person or population?
+How might one apply the FEP to a particular particle, person or population?
 The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing.
@@
 If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed.
-As noted above, technical term for this relationship between real-world system and generative model is “entailment”: where a system that conforms to the dependence structure of the generative model is said to entail that generative model.
+As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
 In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory.
@@
 In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in.
-And even if one could, the internal states would be forever sequestered behind blanket states.
+Even if one could, the internal states would be forever sequestered behind blanket states.
 In short, applications of the FEP place the generative model centre stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
@@
 It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories.
-In developing a Bayesian mechanical account of the dynamics of systems, we need not assume that the particle itself is literally performing inference.
+In developing an inferential account of the dynamics of systems, we need not assume that the particle itself is literally performing inference.
 [^3] Ways of defending stronger, increasingly literal versions of this positions are available as well.
@@
 In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [8, 9]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled.
-This allegation constitutes a fallacy, which one might call the map-territory fallacy fallacy.
+We claim this allegation constitutes a fallacy (one might call this a map-territory fallacy fallacy).
 Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling.
@@
 In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [13].
-The It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
+It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
 Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature.
-In fairness, these ambiguities and idiosyncrasies may lend themselves to misinterpretation.
+In fairness, these ambiguities and idiosyncrasies lend themselves to misinterpretation.
 We now discuss two such cases.

### v3

@@
 This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic.
-The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed.
+The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed.
 We thank the organisers and audiences of both.
 DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York.
+KF and DARS acknowledge support from the VERSES Research Lab.
+Funding declaration
+There is no specific funding to be acknowledged.
 1 Introduction
@@
 After reviewing prior work on the FEP, we identify a particular issue that the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way, offering advantages in terms of tractability and explanation, by using gradients of variational free energy to describe the dynamics of interacting systems.
+One of the points this paper will emphasise is that using this framework does not necessarily presuppose that a system actually performs inference on its environment.
+It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories.
+We need not assume that the particle itself is literally performing inference when we build models under our framework.
+What is at stake is an ‘as if’ description: we simply assume that there exists a quantity that varies systematically with the dynamics of the system; it so happens that surprisal (or, equivalently, variational free energy) is a good choice of quantity for physical reasons, and that the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
+The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane.
+It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems.
+However, it is not necessarily a faithful representation of the system in the literalist sense.
+See also [6].
+To that end, after describing the advantages of this FEP we will then argue that there is no conflation of the modelled system and our scientific model of it—in fact, we describe how they are carefully distinguished.
+This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful.
+We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object.
+That is to say, a model of an object can always be written as if it were modelling or inferring the structure of its environment.
+To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world), and the other, between the generative model and the variational density.
+The structure of the paper is as follows.
+We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients.
+This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer.
+To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts.
+Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [7]; see also [8, 3]).
+More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles).
+This partition induces a statistical boundary called a Markov blanket [9], which captures conditional independence relations amongst subsets of the system [2, 10].
+The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track each others’ statistical properties (dynamically) across the blanket [11].
+Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
+We then argue that there is no ambiguity between model and modelled real-world system in this formalism.
+We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
+Accounts such as [6] have encouraged practitioners of the FEP to be explicit about this point.
+Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems.
+In particular, we do not claim that the objects in the system themselves must literally be performing inference.
+Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
+In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system.
+(This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all.
+In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [12, 13], the FEP is a natural expression of statistical physics as a process of making models.
+More broadly we claim that this echoes the sentiments of Wittgenstein, providing a constraint on modelling in terms of logically necessary preconditions for sense-making and the parsing of sensory streams [7].
+For those already familiar with the FEP, this paper can be thought of as going back to basics, emphasising its foundation in inverting generative models (written as SDEs in generalised coordinates), using a path integral scheme for minimising variational free energy [14].
+2 Summary remarks on the free energy principle
+2.1 Fundamentals
+Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [11, 3, 4] for a more extended look).
+The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [15], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 16, 11].
+Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a minimisation of a variational free energy functional.
+We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state density $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$.
+After Pearl (see [9]) we refer to $B$ as the Markov blanket of the system.
+Typically, $B$ will itself consist of two different kinds of states—sensory states, providing the system signals about its environment, and active states, providing interventions of the system on the environment.
+The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [17].
+The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where ^$\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3].
+Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution.
+In particular: if we assume that $\hat{\eta}_{b}$ parametrises $p(\eta\mid b)$, then we have made $\mu$ into an estimator of the statistics of the environment.
+The FEP is simplest when we have a non-equilibrium steady state density, in which case the following claim holds: if the system remains on some attractor of states, it looks like it minimises variational free energy.
+This is true in the following sense.
+It is possible to express the dynamics of a random variable with a non-equilibrium steady state density in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
+[…]
+Moreover, the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from the most likely state, as well as physically-derived quantities [22, 17, 23], making this a meaningful equation.
+(See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics.)
+Let us make the Laplace assumption such that the most likely state is the expectation. Then our generic SDE (2.1.1) tends towards the fixed point solution
+$\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,max}}\log p(\mu,b).$
+To make use of this first observation, we now postulate a conditional density over external states, $q(\eta;\hat{\eta}_{b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ of the environment.
+Under the Laplace assumption, in a neighbourhood of its peak the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance.
+This guarantees that finding the right value of $\hat{\eta}_{b}$ is sufficient for $q$ to equal the posterior almost surely.
+More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
+Once again, this implies that the system’s average behaviour encodes an inference about its environment, in the sense of parameter estimation.
+Inference by parameter estimation implies the KL divergence between the posterior and the true density vanishes as the parameter varies. We will write
+[…]
+Now observe that since the divergence term is non-negative, $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
+$F(\mu,b)=D_{\mathrm{KL}}(q(\eta;\sigma(\mu))\|\,p(\eta\mid b))-\log p(\mu,b)\geqslant-\log p(\mu,b)$ (2.1.2)
+with equality precisely when the KL term vanishes.
+We conclude that the minimisation of surprisal implies the vanishing of the divergence term.
+Let us apply the Laplace approximation [25] such that the expected state is approximately the mode of the density.
+It follows that the most probable state is the least surprising one.
+Hence, for physical reasons, we expect the surprisal to be minimised.
+From that reasoning, (2.1.1) has the equivalent form
+[…]
+Our reasoning so far implies this suffices as an approximate description of all Markov blanketed systems.
+What we have done is used the fact that the dynamics of a (steady state) system minimise surprisal on average, and argued that whenever equipped with a Markov blanket they also minimise variational free energy; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on variational free energy.
+This gradient flow is what we identify as actually being the process (i.e., the performance) of inference.
+Explicit examples of this can be found in [24, 26, 11].
+The particular partition becomes crucial here, since this is what we leverage in our model to incorporate some unknown quantity (external states beyond the blanket) that the system can do inferences about at all.
+Variational free energy, being a tractable (i.e., easily computed) upper bound on surprisal, already makes this a useful tool to have—however we will go further and claim supposing the existence of a minimised variational free energy, with its interpretation as inferring a target density by approximating it with a variational density, has a use in interpreting models in statistical physics in a new way.
+We will remark that much of this story can be told for the trajectories of a system and distributions over path spaces, but with many modifications [11, 4], none of which are necessary to make the point sought in this paper.
+2.2 Analysis and interpretation
+We have, in the previous subsection, introduced a vanishing KL divergence term to our SDE-based model of a physical system, which we have then stipulated to vanish.
+Though seemingly vacuous, this move opens a world of interesting interpretations for models built according to this reasoning, by connecting the physically-given dynamics of the mode of the system to inference performed by the mode of an estimator.
+For example: by existing a particular way inside the joint state space, the particle can be seen as encoding (via its mode) an inference about a generative model over the entire joint state space, resulting from carving out some ‘niche’ within it.
+That is to say, the environment must likely be some certain way for the system to exist the way it does.
+On the other hand, if free energy (or surprisal) remains irreducibly high, it is likely that the particle occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water).
+Correspondingly, its inference about the generative model would be incorrect, concentrating mass in the wrong areas of the joint state space.
+These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
+Here, variational free energy is a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system.
+The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function.
+Variational free energy arises as a Lyapunov function for the straightforward reason that the most probable state is the least surprising one, and that the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from that most likely state [22, 17, 23]—combined with the inequality in (2.1.2).
+Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy.
+Said another way, the crucial idea is that the dynamics of a system do inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g.
+Freidlin–Wentzell, and noting that the KL divergence is stipulated to be minimised by a near-tautology that the system will be at its average state on average.
+In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution.
+Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like.
+Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence.
+If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
+The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration.
+Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [11]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
+Similarly, note the direction of the implication in our claims.
+In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [16].
+We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [11, §IIB] where this argument is spelt out).
+This happens in the background, so to speak.
+Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing.
+Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation.
+All this is made possible by the existence of a particular partition.
+Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
+However—note that when the Markov blanket consists of sensory states and active states, this inference can immediately be written as part of a control problem [27, 28, 29, 30, 31, 32, 33], lending it an even more sophisticated interpretation.
+From the viewpoint of model building, since internal states are by construction unobservable, we can in principle tell any story that we like about them—provided that this story is supported well enough by observable evidence.
+One such story is that those internal states look as if they parameterise a variational density over external states.
+Taking this explanation seriously, and leveraging it as an approach to modelling physical systems, is applying the FEP.
+2.3 Empirical validation
+In this section we bolster the claim that modelling physical systems with variational free energy is advantageous by reviewing some simulations self-organisation and self-assembly—namely, that of cellular morphogenesis.
+In these simulations we will see that introducing the generative model allows us to model the system self-organising as if it ‘knew its place’ within a larger system—inferring the unknowable external states by the minimisation of variational free energy.
+[^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/).
+These simulations can be reproduced using the morphogenesis feature and the data provided there.
+Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy.
+The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy.
+In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences.
+A group of eight cells was simulated with states consisting of the physical locations on a two-dimensional grid ($\psi_{x}$) of each cell and the cell type of each cell (measured by its cell signalling profile, a four-dimensional vector of signals expressed $\psi_{c}$).
+For cell $i$, $(\psi_{xi},\psi_{ci})$ is an internal state, and the remaining tuples $(\psi_{xj},\psi_{cj})$ constitute external states.
+We assume the active states of each cell couple to $\psi_{c}$ so that the actions the cell can take are manipulation of the signals expressed in the environment.
+We moreover assume the active states fluctuate quickly enough that from the viewpoint of the environment their dynamics are negligible, i.e.
+$\psi\approx a$, constituting an adiabatic approximation
+$\tau\partial_{t}\psi=a-\psi$
+with large time constant $\tau$. Hence the $\psi_{xj}$’s become external states coupled to active states in the Markov blankets of the other cells.
+We will additionally assume an exponential decay in space of extracellular signal concentration as a function of the cell’s own signalling. This means the cell can sense
+$\lambda_{i}(\psi_{x},\psi_{c})=C\sum_{j}\psi_{cj}\exp\big(\absolutevalue{\psi_{xi}-\psi_{xj}}\big{missing}),\quad C\in[0,1]$
+Each cell also has sensory states detecting the cell’s current location and the concentration of chemical signals produced by the cell.
+We will assume these are simply noisy estimates of $(\psi_{xi},\psi_{ci})$ with noise distributed like a Gaussian of precision $\Pi^{(1)}=e^{-16}$, for the full sensory state
+$s_{i}=[\psi_{xi}+\omega,\psi_{ci}+\omega,\lambda_{i}(\psi_{x},\psi_{c})+\omega].$
+A point in the state space $(\psi_{x}^{*},\psi_{c}^{*})$ was designated as a target morphology, giving prescribed locations and cell types of all cells.
+This point consisted of the physical form of a head, body, and tail, and with different cell types in each component of the organism.
+We then equipped each cell with a generative model, specifying it as follows.
+The mean was given in terms of how environmental states are mapped to sensory states,
+$g(\psi)_{i}=\frac{\exp(\psi_{i})}{\sum_{j}\exp(\psi_{j})}\left[\begin{array}[]{c}\psi_{x}^{*}\\ \psi_{c}^{*}\\ \lambda(\psi_{x}^{*},\psi_{c}^{*})\\ \end{array}\right]$
+with the softmax function returning the expected identity of each cell.
+A precision $\Pi^{(2)}=e^{-2}$ was also set over external states.
+A prediction error for how the generative model matched with sensations was given in the form
+$\varepsilon_{i}=[s_{i}-g(\psi)_{i}]$
+The generative model was operationalised by following the procedure in [14].
+A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time.
+A derivative operator $D$ was defined for ordinary differential equations on this space.
+The entire trajectory was given a coordinate in function space
+$\tilde{\psi}=[\psi,\psi^{\prime},\psi^{\prime\prime},\ldots].$
+Treating (2.1.3) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final equations of motion given the generative model were obtained as
+$\begin{split}f_{\tilde{\psi}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{\psi_{i}}-\Gamma_{\psi_{i}})\nabla_{\tilde{\psi}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{\psi}_{i}-\nabla_{\tilde{\psi}_{i}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\psi}_{i}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{a_{i}}-\Gamma_{a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\\ \partial_{t}\tilde{a}_{c}&=D\tilde{a}_{c}-\Pi^{(1)}_{c}\tilde{\varepsilon}_{c}+\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}\\ \partial_{t}\tilde{a}_{x}&=D\tilde{a}_{x}-\nabla_{x}\tilde{s}_{x}\cdot\Pi^{(1)}_{x}\tilde{\varepsilon}_{x}+\nabla_{x}\tilde{s}_{\lambda}\cdot\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}.\end{split}$ (2.3.1)
+We then simulated the dynamics of each cell.
+((a))
+((b))
+((c))
+Figure 1: Plots of the sensed extracellular gradients (1(a)), encoded signal in the cells (1(b)) and softmax expectations of the identities of each cell (1(c)).
+In Figure 1(a) colours denote different signal expressions and black dots denote no data, with the final locations of the cells starred.
+((a)) Migration of cells in time.
+As the cells differentiate they are marked with red, yellow, green, or blue, according to the signal they follow, with the gradient denoting the level of how strongly a signal is expressed.
+((b)) Decrease of free energy in time. A decrease of free energy in time is noted as the cells migrate to the final states corresponding to the sensed target signal.
+Figure 2: Here morphogenesis is demonstrated by cells migrating to align with sensory input, modelled by a concomitant decrease of free energy in time.
+Figure 3: Dynamics of each cell, with final locations starred.
+The gradients here match those of Figure 2(a).
+It can be seen that (speaking somewhat heuristically) the cell moves to fulfil its expectations about the signals it will encounter, whilst expressing the signals associated with its current beliefs about its place in the target ensemble.
+3 Some preliminary philosophical considerations
+We will now enter a more philosophical discussion concerning the use of the FEP, guiding its application and interpretation.
+The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket.
+This means that the FEP is what it says on the tin: it is a principle.
+Like the principles of least action or of maximum entropy, the FEP is a machinery that produces models [3].
+How might one apply the FEP to a particular particle, person or population?
+The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing.
+What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation.
+If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed.
+As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
+If we now move to a meta-theoretical perspective and consider one particle (e.g., a scientist or philosopher) observing another particle, we have the interesting situation where the modeller may impute a particular generative model that best explains the mechanics of the observed particle.
+Is this meta-model a map or a territory?
+[^2] This meta-theoretical move is commonplace in practical applications of the FEP and is sometimes referred to as meta-Bayesian inference or, more simply, observing the observer [35].
+Practically, it involves optimising the parameters of a generative model, such that under the FEP (the ideal Bayesian observer assumption) the observable behaviour of the particle is rendered the most likely.
+This application of the FEP is often described as computational phenotyping [36].
+Using a generative model to capture the formal structure of the environment—and to distinguish it from the model that is carried by particles (the variational density)—can be read as a new take on the nouvelle AI idea that physical systems are their own best models; or as a new take on the good regulator theorem.
+For instance, in work on robotics following the tradition of embodied cognition [37, 38], practitioners have eschewed the construction of agents with internal representations of some environment.
+In these systems, the physics and geometry of the situation were sufficient to endow these agents with the capacity to couple to an environment.
+The world is, on this view, its own best representation.
+In our view, this approach dovetails nicely with FEP-theoretic modelling, where the generative model in FEP-theoretic constructions is just a joint probability density defined over all the states (or paths) of a system.
+In other words, the generative model is just our representation of the world as we believe it to be; and this model need not be encoded directly in the particle or agent.
+Given a generative model or Lagrangian of the appropriate sort, we can show that the system evinces a particular partition (i.e., contains particles); and we can show that subsets of the system track each other, where tracking means inferring or becoming the sufficient statistics of probabilistic beliefs about their external states.
+We interpret this tracking as a form of inference, namely, variational or approximate Bayesian inference under a generative model (i.e., the map).
+In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory.
+Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself.
+In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in.
+Even if one could, the internal states would be forever sequestered behind blanket states.
+Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
+4 A map of that part of the territory that behaves as if it were a map
 Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled).
@@
 By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
-After describing the advantages of this FEP we will then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished.
-We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy.
-This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful.
-We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object.
-That is to say, a model of an object can always be written as if it were modelling or inferring the structure of its environment.
-To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
-To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts.
-Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [15]; see also [13, 16, 3]).
-More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles).
-This partition induces a statistical boundary called a Markov blanket [17], which captures conditional independence relations amongst subsets of the system [2, 18].
-The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket.
-Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
-We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
-Accounts such as [10] have encouraged practitioners of the FEP to be explicit about this point.
-Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems.
-In particular, we do not claim that the objects in the system themselves must literally be performing inference.
-Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
-The structure of the paper is as follows.
-We first present a conceptual introduction to the FEP—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients.
-This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer.
-We then argue that there is no ambiguity between model and modelled real-world system in this formalism.
-We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other.
-In summary we propose that the FEP provides a map of any possible map whatsoever—of, or held by—a physical system.
-(This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all.
-In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [20, 21], the FEP is a natural expression of statistical physics as a process of making models.
-We remark that we will not claim to invalidate the map-territory fallacy itself.
-That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]).
-Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification.
-Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
-2 Summary remarks on the free energy principle
-Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [19, 3, 4] for a more extended look).
-The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [22], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 23, 19].
-Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system.
-The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function.
-In the FEP, this Lyapunov function is surprisal or variational free energy, depending on the setting.
-In other words, to understand the claim that the FEP models a system as doing inference, it is crucial to understanding the ‘inference’ component of surprisal minimisation.
-Surprisal is simply the negative log probability.
-Variational free energy is a tractable (i.e., easily computed) upper bound on surprisal, where minimising variational free energy allows us to infer a target density by approximating it with a variational density.
-The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration.
-Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
-We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state distribution $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$.
-After Pearl [17] we refer to $B$ as the Markov blanket of the system.
-We then postulate a conditional density over external states, $q(\eta;\hat{\eta}_{b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ of the environment.
-For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance.
-This guarantees that finding the right value of $\hat{\eta}_{b}$ is sufficient for $q$ to equal the posterior almost surely.
-More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
-The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics, following e.g.
-Freidlin–Wentzell.
-In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal.
-This gradient flow is what we identify as actually being the process (i.e., the performance) of inference.
-Explicit examples of this can be found in [24, 25, 19].
-More specifically, the FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [23, Lemma 4.3].
-In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$, meaning that for as long as the particle is at state $\hat{\mu}_{b}$, the particle models its environment by estimating the parameters of its distribution.
-Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like.
-Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence.
-If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
-Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy
-[…]
-Of physical interest [^1] See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics.
-is the observation that the divergence term is non-negative, meaning $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
-$F(\mu,b)=D_{\mathrm{KL}}(q(\eta;\sigma(\mu))\|p(\eta\mid b))-\log p(\mu,b)\geqslant-\log p(\mu,b)$
-with equality precisely when the KL term vanishes. From that reasoning, writing
-[…]
-in a neighbourhood of the mode we can equivalently write
-[…]
-Though seemingly vacuous, this move opens a world of interesting interpretations for models built according to this reasoning.
-For example: by existing a particular way inside the joint state space, the particle can be seen as encoding (via its mode) an inference about a generative model over the entire joint state space, resulting from carving out some ‘niche’ within it.
-That is to say, the environment must likely be some certain way for the system to exist the way it does.
-On the other hand, if free energy (or surprisal) remains irreducibly high, it is likely that the particle occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water).
-Correspondingly, its inference about the generative model would be incorrect, concentrating mass in the wrong areas of the joint state space.
-These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
-Similarly, note the direction of the implication in our claims.
-In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [23].
-We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [19, §IIB] where this argument is spelt out).
-This happens in the background, so to speak.
-Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing.
-Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation.
-All this is made possible by the existence of a particular partition.
-Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
-From the viewpoint of model building, since internal states are by construction unobservable, we can in principle tell any story that we like about them—provided that this story is supported well enough by observable evidence.
-One such story is that those internal states look as if they parameterise a variational density over external states.
-Taking this explanation seriously, and leveraging it as an approach to modelling physical systems, is applying the FEP.
-3 Some preliminary philosophical considerations
-We will now enter a more philosophical discussion concerning the use of the FEP.
-The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket.
-This means that the FEP is what it says on the tin: it is a principle.
-In the physics, this means that the FEP is the foundation of a method to be applied [3].
-How might one apply the FEP to a particular particle, person or population?
-The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing.
-What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation.
-If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed.
-As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
-In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory.
-Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself.
-In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in.
-Even if one could, the internal states would be forever sequestered behind blanket states.
-In short, applications of the FEP place the generative model centre stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
-4 A map of that part of the territory that behaves as if it were a map
 We argue that the FEP can be understood, metaphorically, a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map.
@@
 We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
-The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane.
-It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems.
-However, it is not necessarily a faithful representation of the system in the literalist sense.
-See also [10].
 In summary, our scientific, FEP-theoretic model is, metaphorically, a map that allows us to say that some (internal) subset of the system looks as if it possesses a map; which we interpret formally as tracking the statistics of another (external) subset.
@@
 This is an answer to allegations that FEP-theoretic modelling commits the map-territory fallacy: namely, it shows that there is no conflation of the map and the territory, in the sense of the map we postulate as being embodied by the system within the mathematical machinery of the FEP.
-Taking all the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system.
+We remark that we do not claim to invalidate the map-territory fallacy itself.
+That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [43]).
+Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification.
+Instead, we are addressing the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
+Taking all of the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system.
 This echoes seminal work in philosophy by Wittgenstein [51].
@@
 It is noteworthy that the manner in which these constructs are deployed in the FEP-theoretic formulation is markedly different from previous Bayesian approaches.
-This has seldom been noted in the literature (with a few exceptions, e.g., [13]); and even when it is discussed, additional issues tend to undermine the clarity of the exposition (as noted by [9]).
+This has seldom been noted in the literature (with a few exceptions, e.g., [8]).
 In more traditional machine learning approaches, e.g., in the literature on the Helmholtz machine [52], the generative model and the recognition model correspond to the feedforward and the feedback passes that together constitute the process of data-driven Bayesian inference.

### v4

@@
 The ensuing variational free energy principle (FEP) has certain advantages over using random dynamical systems explicitly, notably, by being more tractable and offering a parsimonious explanation of why the joint system evolves in the way that it does, based on the properties of the coupling between system components.
-Using the FEP allows us to model the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics.
-In short, we argue that using generative models to represent and track relations among subsystems leads us to a particular statistical theory of interacting systems.
-Conversely, this theory enables us to construct nested models that respect the known relations among subsystems.
+The FEP is a method whose use allows us to build a model of the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics.
+In short, we argue that using generative models to represent and track relations amongst subsystems leads us to a particular statistical theory of interacting systems.
+Conversely, this theory enables us to construct nested models that respect the known relations amongst subsystems.
 We point out that the fact that a physical object conforms to the FEP does not necessarily imply that this object performs inference in the literal sense; rather, it is a useful explanatory fiction which replaces the ‘explicit’ dynamics of the object with an ‘implicit’ flow on free energy gradients—a fiction that may or may not be entertained by the object itself.
@@
 DARS acknowledges support from the Einstein Chair programme at the Graduate Centre of the City University of New York.
-KF and DARS acknowledge support from the VERSES Research Lab.
+KJF and DARS acknowledge support from the VERSES Research Lab.
 Funding declaration
 There is no specific funding to be acknowledged.
+Author contributions
+The list of authors has been ordered alphabetically by surname.
+The CRediT contributions are as follows: KJF (formal analysis, software, writing—review and editing), MJDR (conceptualisation, writing—review and editing), DARS (conceptualisation, formal analysis, software, visualisation, writing—original draft).
 1 Introduction
@@
 However, it is not necessarily a faithful representation of the system in the literalist sense.
-See also [6].
+(See also [6].)
 To that end, after describing the advantages of this FEP we will then argue that there is no conflation of the modelled system and our scientific model of it—in fact, we describe how they are carefully distinguished.
@@
 The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [17].
-The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where ^$\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3].
+The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3].
 Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution.
@@
 This is true in the following sense.
-It is possible to express the dynamics of a random variable with a non-equilibrium steady state density in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
+It is possible to express the dynamics of a random variable with a non-equilibrium steady state density
+[…]
+in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
+[…]
+where $Q(x)$ is skew-symmetric everywhere, $\Gamma(x)$ is positive (semi-)definite everywhere and satisfies $2\Gamma(x)=D(x)D^{\top}(x)$, and $p^{*}(x)$ is the non-equilibrium steady state density.
+This decomposition holds whenever there exists a $Q(x)$ such that
+$\nabla\times[Q(x)+\Gamma(x)]^{-1}f(x)=0$
 […]
@@
 […]
-Now observe that since the divergence term is non-negative, $F(\mu,b)$ is an upper bound on the surprisal $-\log p(\mu,b)$,
+Now observe that since the divergence term is non-negative, $F(\mu,b)$ is an upper bound on the surprisal $\log p(\mu,b)$,
 $F(\mu,b)=D_{\mathrm{KL}}(q(\eta;\sigma(\mu))\|\,p(\eta\mid b))-\log p(\mu,b)\geqslant-\log p(\mu,b)$ (2.1.2)
@@
 […]
-Our reasoning so far implies this suffices as an approximate description of all Markov blanketed systems.
+Evidently, when the inference gap is minimised, we have
+[…]
+on the nose. Our reasoning so far implies this suffices as an approximate description of all Markov blanketed systems.
 What we have done is used the fact that the dynamics of a (steady state) system minimise surprisal on average, and argued that whenever equipped with a Markov blanket they also minimise variational free energy; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on variational free energy.
@@
 Taking this explanation seriously, and leveraging it as an approach to modelling physical systems, is applying the FEP.
-2.3 Empirical validation
-In this section we bolster the claim that modelling physical systems with variational free energy is advantageous by reviewing some simulations self-organisation and self-assembly—namely, that of cellular morphogenesis.
+3 Empirical validation
+In this section we bolster the claim that modelling physical systems with variational free energy is advantageous by reviewing some simulations of self-organisation and self-assembly.
+Two toy models will be studied—cellular morphogenesis, and an ensemble of excitable cells.
 In these simulations we will see that introducing the generative model allows us to model the system self-organising as if it ‘knew its place’ within a larger system—inferring the unknowable external states by the minimisation of variational free energy.
-[^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/).
-These simulations can be reproduced using the morphogenesis feature and the data provided there.
+3.1 Cellular morphogenesis
 Starting from an ensemble of undifferentiated cells at the same location, we wanted to simulate their migration and differentiation using the minimisation of variational free energy.
@@
 In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences.
+[^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/).
+They can be reproduced using the morphogenesis feature and the data provided there.
 A group of eight cells was simulated with states consisting of the physical locations on a two-dimensional grid ($\psi_{x}$) of each cell and the cell type of each cell (measured by its cell signalling profile, a four-dimensional vector of signals expressed $\psi_{c}$).
-For cell $i$, $(\psi_{xi},\psi_{ci})$ is an internal state, and the remaining tuples $(\psi_{xj},\psi_{cj})$ constitute external states.
-We assume the active states of each cell couple to $\psi_{c}$ so that the actions the cell can take are manipulation of the signals expressed in the environment.
-We moreover assume the active states fluctuate quickly enough that from the viewpoint of the environment their dynamics are negligible, i.e.
+For cell $i$, $(\psi_{xi},\psi_{ci})$ is an internal state $\mu$, and the remaining tuples $(\psi_{xj},\psi_{cj})$ constitute external states $\eta$.
+Each cell was equipped with active states $(a_{x},a_{c})$ corresponding to modulations of its position and chemotactic signalling profile, respectively.
+We assume the active states fluctuate quickly enough that from the viewpoint of the environment their dynamics are negligible, i.e.
 $\psi\approx a$, constituting an adiabatic approximation
 $\tau\partial_{t}\psi=a-\psi$
-with large time constant $\tau$. Hence the $\psi_{xj}$’s become external states coupled to active states in the Markov blankets of the other cells.
+with large time constant $\tau$.
+In this way the cell’s active states become the external states offered to other cells.
+(Note that this assumption is purely for convenience, simplifying the analytic expressions for certain quantities; it is by no means necessary in the following discussion, and the example to follow will not use it.
+What is at stake in terms of the realism of the simulation is: the adiabatic assumption excludes the possibility of modelling cell migration or signal expression as a function of action.)
 We will additionally assume an exponential decay in space of extracellular signal concentration as a function of the cell’s own signalling. This means the cell can sense
-$\lambda_{i}(\psi_{x},\psi_{c})=C\sum_{j}\psi_{cj}\exp\big(\absolutevalue{\psi_{xi}-\psi_{xj}}\big{missing}),\quad C\in[0,1]$
+[…]
 Each cell also has sensory states detecting the cell’s current location and the concentration of chemical signals produced by the cell.
@@
 $s_{i}=[\psi_{xi}+\omega,\psi_{ci}+\omega,\lambda_{i}(\psi_{x},\psi_{c})+\omega].$
+In summary, the active states are (i) where a cell chooses to be and (ii) what a cell chooses to release.
+Those actions are the external states being offered to other cells under the adiabatic assumption, and through the kernel above they superpose into a shared field which every cell samples.
 A point in the state space $(\psi_{x}^{*},\psi_{c}^{*})$ was designated as a target morphology, giving prescribed locations and cell types of all cells.
+This preferred morphology was encoded in a prior and specified desired positions and signal expressions for each clone or cell in terms of the cell’s expectations regarding sensory signals and spatial configurations.
 This point consisted of the physical form of a head, body, and tail, and with different cell types in each component of the organism.
+The target sensation under $\lambda$ is plotted in Figure 1(a).
+In real situations one can imagine a genetic encoding of the target morphology, with epigenetic dynamics being active inference.
+A possible example of this is given in Figure 1(b) where the target is encoded in binary codons for location and expression.
 We then equipped each cell with a generative model, specifying it as follows.
@@
 with the softmax function returning the expected identity of each cell.
+In Figure 1(c) we show the softmax function, interpreted as the posterior beliefs that each cell (column) occupies a particular place in the ensemble (rows).
 A precision $\Pi^{(2)}=e^{-2}$ was also set over external states.
@@
 $\varepsilon_{i}=[s_{i}-g(\psi)_{i}]$
-The generative model was operationalised by following the procedure in [14].
+The model was operationalised by following the procedure in [14].
 A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time.
@@
 $\tilde{\psi}=[\psi,\psi^{\prime},\psi^{\prime\prime},\ldots].$
-Treating (2.1.3) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final equations of motion given the generative model were obtained as
-$\begin{split}f_{\tilde{\psi}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{\psi_{i}}-\Gamma_{\psi_{i}})\nabla_{\tilde{\psi}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{\psi}_{i}-\nabla_{\tilde{\psi}_{i}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\psi}_{i}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})&=(Q_{a_{i}}-\Gamma_{a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\psi}_{i})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\\ \partial_{t}\tilde{a}_{c}&=D\tilde{a}_{c}-\Pi^{(1)}_{c}\tilde{\varepsilon}_{c}+\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}\\ \partial_{t}\tilde{a}_{x}&=D\tilde{a}_{x}-\nabla_{x}\tilde{s}_{x}\cdot\Pi^{(1)}_{x}\tilde{\varepsilon}_{x}+\nabla_{x}\tilde{s}_{\lambda}\cdot\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}.\end{split}$ (2.3.1)
+Now by the decomposition (2.1.1), we have the equations of motion
+[…]
+for each cell.
+[^2] This follows from the marginal flow lemma of [35].
+This is where the interpretation as inference which we supply as modellers comes into play.
+We cannot compute the solution to (3.1.1) since the surprisal term yields an intractable integral.
+However, under the generative model we have equipped our system with, we can instead treat the dynamics of the cell under (3.1.1) as maximising log-evidence for the sensations of the cell, and hence as a variational problem—as posed throughout §2.
+We will introduce the variational distribution
+$),q(\tilde{\psi}\mid\tilde{\mu})=\mathcal{N}\big(\tilde{\mu};-\nabla_{\tilde{\mu}\tilde{\mu}}\log p(\tilde{\mu},\tilde{s},\tilde{a},\tilde{\mu}$$\big)$,
+a Gaussian with expectation being the coordinate of the internal trajectory in function space $\tilde{\mu}$ and variance being the curvature of the surprisal—computed at $\tilde{\psi}=\tilde{\mu}$—and free energy
+$).F(\tilde{a},\tilde{s},\tilde{\psi})=-\log p(\tilde{s},\tilde{a},\tilde{\mu})+D_{\mathrm{KL}}\big(q(\tilde{\psi}\mid\tilde{\mu})\,\|\,p(\tilde{\psi}\mid\tilde{s},\tilde{a},\tilde{\mu}$$\big)$.
+Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)
+$\begin{split}f_{\tilde{\mu}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{\mu\mu}-\Gamma_{\mu\mu})\nabla_{\tilde{\mu}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{\mu}-\nabla_{\tilde{\mu}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\mu}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{a_{i}a_{i}}-\Gamma_{a_{i}a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\end{split}$ (3.1.2)
+with the following explicit equations for (positional and signalling, respectively) active states
+$\partial_{t}\tilde{a}_{x}$ $=D\tilde{a}_{x}-\nabla_{x}\tilde{s}_{x}\cdot\Pi^{(1)}_{x}\tilde{\varepsilon}_{x}+\nabla_{x}\tilde{s}_{\lambda}\cdot\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}$
+$\partial_{t}\tilde{a}_{c}$ $=D\tilde{a}_{c}-\Pi^{(1)}_{c}\tilde{\varepsilon}_{c}+\Pi^{(1)}_{\lambda}\tilde{\varepsilon}_{\lambda}.$
+As before, if the inference gap is minimised, for instance, in the neighbourhood of the mode, then (3.1.2) is equal to (3.1.1) on the nose.
 We then simulated the dynamics of each cell.
+In Figure 2 we have shown the expected positions and releases of each cell over time, extracted from the solution to (3.1.2), and the decrease in free energy as the cells evolve towards the target state.
+In summary, the cell carries a generative model that predicts what signals it should sense if it were the ‘correct’ type of cell at the ‘correct’ place in the target morphology.
+Those predictions are built from the target locations and target signal combinations, and the model maps internal beliefs about identity through the generative model to predicted sensory consequences.
+The final outcome of this process is shown in Figure 3.
 ((a))
@@
 ((c))
-Figure 1: Plots of the sensed extracellular gradients (1(a)), encoded signal in the cells (1(b)) and softmax expectations of the identities of each cell (1(c)).
+Figure 1: Plots of the target extracellular gradients (1(a)), encoding of the target signal in the cells (1(b)) and softmax expectations of the identities of each cell (1(c)).
 In Figure 1(a) colours denote different signal expressions and black dots denote no data, with the final locations of the cells starred.
@@
 ((b)) Decrease of free energy in time. A decrease of free energy in time is noted as the cells migrate to the final states corresponding to the sensed target signal.
-Figure 2: Here morphogenesis is demonstrated by cells migrating to align with sensory input, modelled by a concomitant decrease of free energy in time.
+Figure 2: Here morphogenesis is demonstrated by cells migrating to align with expectations based on sensory input, modelled by a concomitant decrease of free energy in time.
 Figure 3: Dynamics of each cell, with final locations starred.
 The gradients here match those of Figure 2(a).
-It can be seen that (speaking somewhat heuristically) the cell moves to fulfil its expectations about the signals it will encounter, whilst expressing the signals associated with its current beliefs about its place in the target ensemble.
+It can be seen that (speaking somewhat heuristically) each cell moves to fulfil its expectations about the signals it should encounter, whilst expressing the signals associated with its current beliefs about its place in the target ensemble.
+3.2 Periodically-firing cells
+A similar but more complicated example is a ring of excitable, gap-junction-coupled cells whose generative model encodes a periodic target waveform.
+Each cell’s internal states infer the current phase of that cycle from its membrane sensors, whilst active states like ion-channel gating act to make sensed signals match the predicted periodic trajectory.
+This model is more expressive because the prior over hidden states is oscillatory, such that the free energy minimum is a limit cycle: the population converges to a stable circulating orbit in state space rather than to a fixed point.
+We will define external $\eta_{t}$ as a stimulus in the world, $s_{t}$ as a sensory state, $\mu_{t}$ as the internal state estimating the phase of the cell, and $a_{t}$ as the state of some actuator or motor acting on the environment to make observations conform to predictions.
+For simplicty we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e.
+that the likelihood satisfies
+[…]
+at steady state. We will also introduce a likelihood over outcomes of actions
+[…]
+in place of the simplifying adiabatic assumption used in the previous example.
+The system’s predictions of sensations given a phase is the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
+[…]
+There will finally be a uniform prior on $\mu$ as a circle-valued variable.
+In total we have a one-step generative model (conditioned on the internal phase):
+$s\mid\eta\sim\mathcal{N}(\eta,\sigma_{s}^{2}),\qquad\eta\mid a\sim\mathcal{N}(a,\sigma_{\eta}^{2}),\qquad a\mid\mu\sim\mathcal{N}(A\cos\mu,\sigma_{a}^{2}).$
+The generative model is linear in $(s,\eta,a)$ and mildly nonlinear in $\mu$ through the cosine function.
+To keep this example brief—and to make the surprisal analytically soluble, such that some direct comparisons can be made between (2.1.1) and (2.1.3)—we will not create any hidden states to make inferences about beyond the environmental state, so our generative model is not hierarchical and generalised coordinates of motion need not be used.
+In that case, using conditional independence, we can write at steady state (up to additive constants)
+$\log p(x)=\frac{(s-\eta)^{2}}{2\sigma_{s}^{2}}+\frac{(\eta-a)^{2}}{2\sigma_{\eta}^{2}}+\frac{(a-A\cos\mu)^{2}}{2\sigma_{a}^{2}}.$
+Now specify a parametrisation of $Q$ and $\Gamma$,
+$Q=\begin{bmatrix}0&q_{\mu s}&q_{\mu a}&0\\ -q_{\mu s}&0&0&0\\ -q_{\mu a}&0&0&q_{a\eta}\\ 0&0&-q_{a\eta}&0\end{bmatrix},\qquad\qquad\Gamma=\mathrm{diag}(\gamma_{\mu},\gamma_{s},\gamma_{a},\gamma_{\eta}).$
+Notice there are no couplings in the flow over the state space between $\mu$ and $\eta$. This is an indication of the Markov blanket.
+Since the generative model is somewhat simple, it is straightforward to write the equations of motion:
+$\begin{split}\partial_{t}\mu_{t}&=q_{\mu s}\partial_{s}L+q_{\mu a}\partial_{a}L-\gamma_{\mu}\partial_{\mu}L+\xi_{t}^{\mu}\\ &=q_{\mu s}\frac{s-\eta}{\sigma_{s}^{2}}+q_{\mu a}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)-\gamma_{\mu}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)+\xi_{t}^{\mu}\\ \partial_{t}s_{t}&=-q_{\mu s}\partial_{\mu}L-\gamma_{s}\partial_{s}L+\xi_{t}^{s}\\ &=-q_{\mu s}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)-\gamma_{s}\frac{s-\eta}{\sigma_{s}^{2}}+\xi_{t}^{s}\\ \partial_{t}a_{t}&=-q_{\mu a}\partial_{\mu}L+q_{a\eta}\partial_{\eta}L-\gamma_{a}\partial_{a}L+\xi_{t}^{a}\\ &=-q_{\mu a}\frac{A\sin\mu}{\sigma^{2}_{a}}(a-A\cos\mu)+q_{a\eta}\left(\frac{\eta-s}{\sigma^{2}_{s}}+\frac{\eta-a}{\sigma_{\eta}^{2}}\right)-\gamma_{a}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)+\xi_{t}^{a}\\ \partial_{t}\eta_{t}&=-q_{a\eta}\partial_{a}L-\gamma_{\eta}\partial_{\eta}L+\xi_{t}^{\eta}\\ &=-q_{a\eta}\left(\frac{a-\eta}{\sigma_{\eta}^{2}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)-\gamma_{\eta}\left(\frac{\eta-s}{\sigma^{2}_{s}}+\frac{\eta-a}{\sigma_{\eta}^{2}}\right)+\xi_{t}^{\eta}\end{split}$ (3.2.1)
+where $\xi^{i}_{t}$ are independent Brownian motions. Notice the solenoidal couplings between sensory states and environmental states preclude any adiabatic assumption.
+Using properties of marginals of jointly Gaussian random variables we can compute the steady state
+[…]
+i.e.
+i.e.
+a uniform phase prior, a Gaussian action prior around $A\cos\mu$, and a sensor likelihood concentrated around the action with variance $\sigma_{s}^{2}+\sigma_{\eta}^{2}$.
+The interpretation of this density is that the triple $(\mu,s,a)$ forms a ribbon-shaped non-equilibrium steady state in $S^{1}\times\mathbf{R}^{2}$ aligned with the subspace $\{a=A\cos\mu,s=a\}$.
+The solenoidal component makes the probability circulate along this ribbon; however, the shape of the ribbon is fixed by the generative terms specified above.
+The system can treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
+[…]
+Completing the square in $\eta$ yields $p(\eta\mid s,a)=\mathcal{N}(m_{\mathrm{post}},v_{\mathrm{post}})$ where
+$v_{\mathrm{post}}=\left(\frac{1}{\sigma_{s}^{2}}+\frac{1}{\sigma_{\eta}^{2}}\right)^{-1},\qquad m_{\mathrm{post}}=v_{\mathrm{post}}\left(\frac{s}{\sigma^{2}_{s}}+\frac{a}{\sigma^{2}_{\eta}}\right)$
+so that the posterior is parametrised by a variance of twice the inverse precision.
+Now let the internal states parametrise a Gaussian posterior expressing beliefs about the world, $q(\eta;\mu,\tau^{2})=\mathcal{N}(\eta;A\cos\mu,\tau^{2})$.
+The free energy can be expressed analytically as
+$F(s,a,\mu)$ $=\operatorname{\mathbb{E}}_{q}[\log q(\eta;\mu,\tau^{2})]-\operatorname{\mathbb{E}}_{q}[\log p(\eta\mid s,a)]-\log p(s,a,\mu)$
+$=-\frac{1}{2}\log 2\pi\tau^{2}-\frac{1}{2}+\frac{1}{2}\log 2\pi v_{\mathrm{post}}+\frac{1}{2v_{\mathrm{post}}}\left(\tau^{2}+(A\cos\mu-m_{\mathrm{post}})^{2}\right)-\log p(s,a,\mu)$
+or, collecting like terms,
+$\frac{1}{2}\left(\log\frac{v_{\mathrm{post}}}{\tau^{2}}+\frac{\tau^{2}+(A\cos\mu-m_{\mathrm{post}})^{2}}{v_{\mathrm{post}}}-1\right)-\log p(s,a,\mu)$
+so that when $\tau^{2}=v_{\mathrm{post}}$ and $m_{\mathrm{post}}=A\cos\mu$, the KL term vanishes.
+Taking gradients of the free energy yields the equations of motion
+$\begin{split}\partial_{t}\mu_{t}&=q_{\mu s}\partial_{s}F+q_{\mu a}\partial_{a}F-\gamma_{\mu}\partial_{\mu}F+\xi_{t}^{\mu}\\ &=q_{\mu s}\frac{s-A\cos\mu}{\sigma_{s}^{2}}+q_{\mu a}(a-A\cos\mu)\left(\frac{1}{\sigma_{\eta}^{2}}+\frac{1}{\sigma_{a}^{2}}\right)\\ &\phantom{--}-\gamma_{\mu}A\sin\mu\left(\frac{s-A\cos\mu}{\sigma_{s}^{2}}+\frac{a-A\cos\mu}{\sigma^{2}_{\eta}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)+\xi_{t}^{\mu}\\ \partial_{t}a_{t}&=-q_{\mu a}\partial_{\mu}F-\gamma_{a}\partial_{a}F+\xi_{t}^{a}\\ &=-q_{\mu a}A\sin\mu\left(\frac{s-A\cos\mu}{\sigma_{s}^{2}}+\frac{a-A\cos\mu}{\sigma^{2}_{\eta}}+\frac{a-A\cos\mu}{\sigma_{a}^{2}}\right)\\ &\phantom{--}-\gamma_{a}(a-A\cos\mu)\left(\frac{1}{\sigma_{\eta}^{2}}+\frac{1}{\sigma_{a}^{2}}\right)+\xi_{t}^{a}.\end{split}$ (3.2.2)
+We then simulated the system as a Kalman update for a linear Gaussian state space channel $\eta\to s$ with input aa.
+[^3] The python code for this simulation is available at https://github.com/DARSakthi/periodically_firing_cells/.
+In practice, this means when simulating the system forwards, we set the parameters of the variational density to those of the posterior such that the variational density $q$ became the (instantaneous) Kalman posterior over the latent world state $\eta$ and the instantaneous precision-weighted prediction errors drove the gradient descent on $F$.
+In Figure 4 sampled trajectories of $\mu_{t}$ under (3.2.1) and (3.2.2) are compared (Figure 5, $a_{t}$, respectively).
+Close correlation is observed between the two flow regimes.
+A corresponding tendency towards zero of the sample mean of the KL term is observed in Figure 6.
+Plotted per-timepoint sample means of the free energy under both flow regimes are contained in Figure 7.
+By adding the precision-weighted prediction error (KL) term to the internal updates, $F$ has a faster decrease and a greater magnitude of decrease under (3.2.2); this is evidently from the prediction error driving the flow of $\partial_{\mu}F$ to align $q(\eta;\mu)$ with $p(\eta\mid s,a)$.
+Figure 4: Evolution of the internal state under $\mathcal{L}$ and under $F$.
+The dynamics of internal states in two regimes, plotted on a polar spiral showing the evolution in the phase in time.
+The dynamics under a gradient descent on $\mathcal{L}$ and on $F$ show excellent agreement over all timescales.
+Figure 5: Evolution of the active state under $\mathcal{L}$ and under $F$. The dynamics of the active state in two regimes were also plotted; again, excellent agreement is observed.
+Figure 6: Decrease of the KL divergence under the flow on $F$. The sample mean ($N=200$) of the KL divergence along trajectories of (3.2.2) is shown.
+Figure 7: Decrease of the free energy under trajectories in both flow regimes.
+The sample mean ($N=200$) of the variational free energy along trajectories of (3.2.1) and along trajectories of (3.2.2) are compared.
+Both drive the free energy towards zero, however, the flow under $F$ decreases more rapidly and attains a lesser minimum closer to zero.
+The reason an inferential interpretation is useful here is in the coding of stimuli in neurones and neuronal ensembles.
+Minimising variational free energy, and the resultant model of internal and active states behaving as if they were performing Bayesian inference, endows the phase of the oscillation with the interpretation of a posterior belief about expected incoming signals in the environment and the Kalman-like updates as modulations of gain to minimise precision-weighted prediction error, rather than merely a position on a limit cycle.
+Under the relationship to stochastic thermodynamics forged in e.g.
+[24], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.
 4 Some preliminary philosophical considerations
@@
 Like the principles of least action or of maximum entropy, the FEP is a machinery that produces models [3].
-How might one apply the FEP to a particular particle, person or population?
+Throughout, ‘as if’ marks an interpretive equivalence: internal flows that minimise variational free energy can be read as variational inference without presupposing explicit, propositional models inside the system.
+Where specific mechanisms are identified (e.g., precision-weighted message passing), this reading coincides with literal implementation; otherwise, it remains an explanatory stance.
+How might one apply the FEP to make a model of a particular particle, person or population?
 The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing.
@@
 We interpret this tracking as a form of inference, namely, variational or approximate Bayesian inference under a generative model (i.e., the map).
+In the context of this meta-question, one may ask how an observer could come to consciously think an observed system is manipulating a model if the observer’s own brain is simply flowing down a free-energy gradient.
+Here, an observer’s conscious report that ‘X is using an internal model’ may arise from some higher-order generative model over another agent’s (or one’s own) internal states, i.e., beliefs about the other system’s blanket-to-internal mappings (what we have called ‘observing the observer’).
+Conscious thought in this context is a higher-level subset of those internal states which summarises beliefs about others (and the self) within a hierarchical model.
+Mechanistically, this report is implemented by the observer’s own free-energy-minimising dynamics.
 In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory.
@@
 Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
+Here both observed and observer are described at the mechanistic level by flows on free energy flows; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.
 5 A map of that part of the territory that behaves as if it were a map
@@
 It is an explicitly inferential representation of the manner in which subsets of a system with a particular partition track each other: in which case, we are describing a part of the territory that, as seen in our map, also behaves as if it were a map.
-5 Some remarks on terminological idiosyncrasies of the FEP literature
-In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [41, 42]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled.
-We claim this allegation constitutes a fallacy (one might call this a map-territory fallacy fallacy).
-Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling.
-Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not.
-In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [8].
-It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
-Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature.
-In fairness, these ambiguities and idiosyncrasies lend themselves to misinterpretation.
-We now discuss two such cases.
-5.1 Generative models and variational densities versus recognition densities
-The first concerns the ambiguous discussion or description of two core constructs in the FEP literature—namely, the generative model and the variational (or recognition) density.
-It is noteworthy that the manner in which these constructs are deployed in the FEP-theoretic formulation is markedly different from previous Bayesian approaches.
-This has seldom been noted in the literature (with a few exceptions, e.g., [8]).
-In more traditional machine learning approaches, e.g., in the literature on the Helmholtz machine [52], the generative model and the recognition model correspond to the feedforward and the feedback passes that together constitute the process of data-driven Bayesian inference.
-In these settings, both the generative model and the recognition model are straightforwardly encoded in the connection weights of the Helmholtz machine: the generative model is encoded in the weighting of the backwards connections, which establish an information channel from high-level latent spaces of the neural network to its sensory layers; whereas the recognition model is encoded in the feedforward connections.
-In this setting, one assumes that any system that implements this pair of functionally differentiated weightings (e.g., a brain) implements a generative and recognition model pair.
-Much of the literature on the FEP is based on this reading of the generative and recognition models: it is often said within this literature that FEP-theoretic generative models are encoded by the brain, much like these would be in a Helmholtz machine.
-However, whilst generative and variational models are also used in the FEP-theoretic formulation, they are deployed in a markedly different way, which we discussed extensively above.
-Overlooking these differences may have contributed to the impression that theorists of the FEP conflate the metaphorical map and the territory.
-Recall that a generative model, generically, is just a joint probability density defined over the states (or paths) of a system.
-In other words, it is our map or representation of the system being modelled and its mechanics; and more particularly, it is our model of the dependencies between the variables of the system, providing us with equations of motion for the whole system (i.e., it is defined over internal, blanket, and external states/paths).
-Technically, we say that the system that we are modelling entails a generative model; whereas the internal states of the particle (look, to the modeller, as if they) estimate the parameters of a conditional density over external states.
-This difference is crucial to understanding what is at stake in FEP-theoretic modelling: we say that one subset of variables over which the generative model is defined (i.e., the internal states) encodes a variational or recognition model over another subset (the external states).
-Entailment is the property that, metaphorically, “the part of the territory that behaves like a map” (the internal states that encode the parameters of the variational density) is itself part of the territory described by our map (the generative model) and so is subject to the mechanics of that territory, as described by our map.
-Entailment means that the relationships that hold between variables, as we represent in the generative model, are obtained as a description of the dynamics of the real-world system.
-The subtle but key point is that the generative model is not encoded by the internal states of the particle being considered: it is not some kind of internal model.
-Rather, the generative model is a description of the dependence relations that obtain in the dynamics of a system: i.e., such-and-such dynamics entail such-and-such equations of motion and dependencies.
-FEP-theoretic modelling does not assume that the generative model is encoded in the brain.
-Arguably, to claim this would indeed constitute a map-territory fallacy.
-Under the FEP, parts of the system itself (i.e., by internal states) look to an external observer as if, or can be modelled as if, they encode the parameters of a variational model.
-However, this raises a question: if it is our scientific model, how can the generative model play a role in inference if only a conditional posterior model is encoded by the system?
-To claim that a system uses or leverages its generative model is, in some sense, a shorthand expressing the fact that the system leverages its own causal structure, and the data patterns that it elicits through action, to generate patterns of adaptive behaviour—all of which are subject to the regularities that are captured in the generative model.
-This is a statement of embodied and enactive cognition [53, 8].
-The system, in having sensory states, also has access to the causal structure in the generative model.
-So, the generative model is our model or map, but it denotes the causal structure of the system being modelled (i.e., the territory).
-This causal structure—and in particular, the presence of a Markov blanket or particular partition—means that some parts of the system (the internal states) encode a variational density over some other parts of the system (the external states); that is, they “act like a map” or model the statistics of external states.
-5.2 Generative models and generative processes
-The second ambiguity concerns the generative model and the so-called “generative process”.
-Some presentations of the FEP sometimes define the generative model as a scientific model of some data-generating generative process; see, for instance, [54, 8]).
-On this reading, the generative model is our “map” of some real-world process, namely, the generative process—which would then correspond to the metaphorical territory.
-In practice, however, in FEP-theoretic modelling, this distinction does not map onto that between a scientific model and real-world system, i.e., onto the distinction between the map and territory.
-Perhaps confusingly, the generative process does not denote (the structure of) the real data-generating process that we are trying to model.
-In FEP-theoretic formulations, the generative process is also part of our scientific model.
-It should be noted that, practically, the distinction between generative model and generative process only comes into play when we are implementing the dynamics of a specific system in simulations.
-When we write code to simulate the system of interest, typically, we need to specify (i) the dependencies that hold between the variables of a system, which we call the generative model; and (ii) the equations of motion that govern the time evolution of the external states per se in our simulation.
-The latter is labelled the ‘generative process’, since it generates the dynamics of external states, which are tracked by internal states.
-The key thing to note, however, is that generative process is also part of our scientific model—it harnesses what we know about the physics of the situation described by the generative model, in which the thing is embedded and ‘lives’.
 6 Conclusion

### v5

@@
 Acknowledgements
-This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic.
+This paper more fully develops the theory outlined in a previous manuscript [mtff], and we are grateful to those listed there for their contributions to preliminary discussions on the topic.
 The account here is largely based on a series of talks given by the third named author at the Department of Philosophy of the University of Vienna, the Laufer Center for Physical and Quantitative Biology, the 2024 Conference on Complex Systems, and the informal discussions that followed.
@@
 In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt.
-One approach to making such models, using variational (approximate Bayesian) inference [2, 3, 4], describes how a physical object reflects the properties of its environment and vice versa.
+One approach to making such models, using variational (approximate Bayesian) inference [Friston2019, ramstead2022bayesian, friston2023path], describes how a physical object reflects the properties of its environment and vice versa.
 In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena.
-More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [5]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action.
+More specifically, we argue that generative modelling—premised on the use of variational inference to model the causes of the signals that a coupled system is receiving [beal2003variational]—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action.
 We use the ensuing variational free energy principle to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling.
@@
 However, it is not necessarily a faithful representation of the system in the literalist sense.
-(See also [6].)
+(See also [Andrews2021].)
 To that end, after describing the advantages of this FEP we will then argue that there is no conflation of the modelled system and our scientific model of it—in fact, we describe how they are carefully distinguished.
@@
 To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts.
-Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [7]; see also [8, 3]).
-More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles).
-This partition induces a statistical boundary called a Markov blanket [9], which captures conditional independence relations amongst subsets of the system [2, 10].
-The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track each others’ statistical properties (dynamically) across the blanket [11].
+Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [Ramstead2020semantics]; see also [Ramstead2019enactive, ramstead2022bayesian]).
+More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so-called ‘particular’ partition, i.e., into objects or particles).
+This partition induces a statistical boundary called a Markov blanket [Pearl1998], which captures conditional independence relations amongst subsets of the system [Friston2019, sakthivadivel2022d].
+The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track each others’ statistical properties (dynamically) across the blanket [classical-physics].
 Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is a form of Bayesian inference.
@@
 We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
-Accounts such as [6] have encouraged practitioners of the FEP to be explicit about this point.
+Accounts such as [Andrews2021] have encouraged practitioners of the FEP to be explicit about this point.
 Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems.
@@
 (This includes the use of the FEP itself by a modeller.) We suggest that this leads to constraints on all possible maps, or models of physical things, arising from what it means to be a map or model of a physical process at all.
-In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [12, 13], the FEP is a natural expression of statistical physics as a process of making models.
-More broadly we claim that this echoes the sentiments of Wittgenstein, providing a constraint on modelling in terms of logically necessary preconditions for sense-making and the parsing of sensory streams [7].
-For those already familiar with the FEP, this paper can be thought of as going back to basics, emphasising its foundation in inverting generative models (written as SDEs in generalised coordinates), using a path integral scheme for minimising variational free energy [14].
+In the fashion of Jaynes, who proposed that doing statistical mechanics is doing inference about the probabilities of microstates given knowledge of macrostates [jaynes1, jaynes2], the FEP is a natural expression of statistical physics as a process of making models.
+More broadly we claim that this echoes the sentiments of Wittgenstein, providing a constraint on modelling in terms of logically necessary preconditions for sense-making and the parsing of sensory streams [Ramstead2020semantics].
+For those already familiar with the FEP, this paper can be thought of as going back to basics, emphasising its foundation in inverting generative models (written as stochastic differential equations (SDEs) in generalised coordinates), using a path integral scheme for minimising variational free energy [friston2010generalised].
 2 Summary remarks on the free energy principle
 2.1 Fundamentals
-Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [11, 3, 4] for a more extended look).
-The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [15], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [3, 16, 11].
+Here we will describe the proposed free energy principle (also directing the reader to prior literature such as [classical-physics, ramstead2022bayesian, friston2023path] for a more extended look).
+The FEP aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference) [DaCosta2021], resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled [ramstead2022bayesian, Sakthivadivel2022b, classical-physics].
 Formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a minimisation of a variational free energy functional.
 We consider a particle (also referred to as a ‘system’) with states $\mu$ coupled to another system with states $\eta$ (henceforth, the ‘environment’) through a set of variables $B$ with values $b$, and a steady state density $p(\eta,b,\mu)$ such that $\mu$ is conditionally independent of $\eta$ given $B$.
-After Pearl (see [9]) we refer to $B$ as the Markov blanket of the system.
+After Pearl (see [Pearl1998]) we refer to $B$ as the Markov blanket of the system.
 Typically, $B$ will itself consist of two different kinds of states—sensory states, providing the system signals about its environment, and active states, providing interventions of the system on the environment.
-The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [17].
-The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [16, Lemma 4.3].
+The coupling at play puts us in a non-equilibrium situation, where a system can be driven by fluxes of energy or matter from other systems in its environment [seifert2012stochastic].
+The FEP begins by saying that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary, and in particular, such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, where $\hat{\,\cdot\,}_{b}$ denotes the conditional expectation given the blanket state $b$ [Sakthivadivel2022b, Lemma 4.3].
 Following that, the aim of the FEP is to understand what the generative model $p(\eta,b,\mu)$ says about the dynamics of the subsystem with states $\mu$, beyond just its marginal distribution.
@@
 […]
-in terms of a gradient flow on the surprisal weighted by a particular matrix field [18, 19, 20, 21], having the form
+in terms of a gradient flow on the surprisal weighted by a particular matrix field [ao2004potential, ma2015, yuan2017sde, da2023entropy], having the form
 […]
@@
 […]
-Moreover, the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from the most likely state, as well as physically-derived quantities [22, 17, 23], making this a meaningful equation.
-(See [24] where this is applied to reproduce certain aspects of stochastic thermodynamics.)
+Moreover, the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from the most likely state, as well as physically-derived quantities [freidlin_wentzell_2012, seifert2012stochastic, seifert2019stochastic], making this a meaningful equation.
+(See [parr2020markov] where this is applied to reproduce certain aspects of stochastic thermodynamics.)
 Let us make the Laplace assumption such that the most likely state is the expectation. Then our generic SDE (2.1.1) tends towards the fixed point solution
@@
 Once again, this implies that the system’s average behaviour encodes an inference about its environment, in the sense of parameter estimation.
-Inference by parameter estimation implies the KL divergence between the posterior and the true density vanishes as the parameter varies. We will write
+Inference by parameter estimation implies the Kullback–Leibler (KL) divergence between the posterior and the true density vanishes as the parameter varies. We will write
 […]
@@
 We conclude that the minimisation of surprisal implies the vanishing of the divergence term.
-Let us apply the Laplace approximation [25] such that the expected state is approximately the mode of the density.
+Let us apply the Laplace approximation [friston2007variational] such that the expected state is approximately the mode of the density.
 It follows that the most probable state is the least surprising one.
@@
 This gradient flow is what we identify as actually being the process (i.e., the performance) of inference.
-Explicit examples of this can be found in [24, 26, 11].
+Explicit examples of this can be found in [parr2020markov, ueltzhoffer2021drive, classical-physics].
 The particular partition becomes crucial here, since this is what we leverage in our model to incorporate some unknown quantity (external states beyond the blanket) that the system can do inferences about at all.
 Variational free energy, being a tractable (i.e., easily computed) upper bound on surprisal, already makes this a useful tool to have—however we will go further and claim supposing the existence of a minimised variational free energy, with its interpretation as inferring a target density by approximating it with a variational density, has a use in interpreting models in statistical physics in a new way.
-We will remark that much of this story can be told for the trajectories of a system and distributions over path spaces, but with many modifications [11, 4], none of which are necessary to make the point sought in this paper.
+We will remark that much of this story can be told for the trajectories of a system and distributions over path spaces, but with many modifications [classical-physics, friston2023path], none of which are necessary to make the point sought in this paper.
 2.2 Analysis and interpretation
@@
 The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function.
-Variational free energy arises as a Lyapunov function for the straightforward reason that the most probable state is the least surprising one, and that the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from that most likely state [22, 17, 23]—combined with the inequality in (2.1.2).
+Variational free energy arises as a Lyapunov function for the straightforward reason that the most probable state is the least surprising one, and that the log-probability is able to describe the probability of all fluctuations of the system in terms of their distance from that most likely state [freidlin_wentzell_2012, seifert2012stochastic, seifert2019stochastic]—combined with the inequality in (2.1.2).
 Since the dynamics towards the mode induce a gradient flow on a divergence between the estimated parametric distribution and the true distribution, the FEP moreover proposes that these dynamics (in virtue of the coupling) can be modelled as a variational inference process, following a gradient flow on the variational free energy.
@@
 The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration.
-Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [11]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
+Minimising surprisal implies that the variational free energy is minimised, (i.e., the object-as-model is a good estimator of its environment [classical-physics]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
 Similarly, note the direction of the implication in our claims.
-In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [16].
-We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [11, §IIB] where this argument is spelt out).
+In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [Sakthivadivel2022b].
+We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [classical-physics, §IIB] where this argument is spelt out).
 This happens in the background, so to speak.
@@
 Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
-However—note that when the Markov blanket consists of sensory states and active states, this inference can immediately be written as part of a control problem [27, 28, 29, 30, 31, 32, 33], lending it an even more sophisticated interpretation.
+However—note that when the Markov blanket consists of sensory states and active states, this inference can immediately be written as part of a control problem [friston2012active, pezzulo2015active, pio2016active, pezzulo2018hierarchical, baltieri2019pid, millidge2020relationship, baioumy2021active], lending it an even more sophisticated interpretation.
 From the viewpoint of model building, since internal states are by construction unobservable, we can in principle tell any story that we like about them—provided that this story is supported well enough by observable evidence.
@@
 The simulation here exploits the fact one can express a Markov blanketed system’s marginal Lyapunov function in terms of a variational free energy function, such that its attracting set can be prescribed in terms of the generative model that defines the free energy.
-In essence, the systems dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences.
-[^1] These simulations were based on [34] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/).
+In essence, the system’s dynamics rest upon each cell inferring its unique identity (in relation to all others) and behaving in accord with those inferences.
+[^1] These simulations were based on [friston2015knowing] and use the DEM toolbox in the MATLAB software SPM12 (https://www.fil.ion.ucl.ac.uk/spm/).
 They can be reproduced using the morphogenesis feature and the data provided there.
@@
 $\varepsilon_{i}=[s_{i}-g(\psi)_{i}]$
-The model was operationalised by following the procedure in [14].
+The model was operationalised by following the procedure in [friston2010generalised].
 A Taylor expansion of trajectories was taken and the statistics of each order of approximation $(\psi,\psi^{\prime},\psi^{\prime\prime},\ldots)$ were modelled at each point in time.
@@
 for each cell.
-[^2] This follows from the marginal flow lemma of [35].
+[^2] This follows from the marginal flow lemma of [Friston2019].
 This is where the interpretation as inference which we supply as modellers comes into play.
@@
 $).F(\tilde{a},\tilde{s},\tilde{\psi})=-\log p(\tilde{s},\tilde{a},\tilde{\mu})+D_{\mathrm{KL}}\big(q(\tilde{\psi}\mid\tilde{\mu})\,\|\,p(\tilde{\psi}\mid\tilde{s},\tilde{a},\tilde{\mu}$$\big)$.
-Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [14, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)
+Treating (3.1.1) as a Bayesian filter of trajectories, from the derivation outlined in [friston2010generalised, equation 3.6] the final drift terms for the equations of motion for the $i$-th cell given the generative model were obtained as (again denoting the internal state $\psi_{i}$ as $\mu$)
 $\begin{split}f_{\tilde{\mu}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{\mu\mu}-\Gamma_{\mu\mu})\nabla_{\tilde{\mu}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{\mu}-\nabla_{\tilde{\mu}}\tilde{\varepsilon}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}-\Pi^{(2)}\tilde{\mu}\\ f_{\tilde{a}_{i}}(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})&=(Q_{a_{i}a_{i}}-\Gamma_{a_{i}a_{i}})\nabla_{\tilde{a}_{i}}F(\tilde{a}_{i},\tilde{s}_{i},\tilde{\mu})=D\tilde{a}_{i}-\nabla_{\tilde{a}_{i}}\tilde{s}_{i}\cdot\Pi^{(1)}\tilde{\varepsilon}_{i}\end{split}$ (3.1.2)
@@
 We will define external $\eta_{t}$ as a stimulus in the world, $s_{t}$ as a sensory state, $\mu_{t}$ as the internal state estimating the phase of the cell, and $a_{t}$ as the state of some actuator or motor acting on the environment to make observations conform to predictions.
-For simplicty we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e.
+For simplicity we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e.
 that the likelihood satisfies
@@
 in place of the simplifying adiabatic assumption used in the previous example.
-The system’s predictions of sensations given a phase is the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
+The system’s predictions of sensations given a phase follow the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
 […]
@@
 The solenoidal component makes the probability circulate along this ribbon; however, the shape of the ribbon is fixed by the generative terms specified above.
-The system can treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
+The system can be treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
 […]
@@
 Under the relationship to stochastic thermodynamics forged in e.g.
-[24], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.
+[parr2020markov], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.
+3.3 Looking ahead
+It is important to consider that we have not quite taken an agnostic approach, i.e., taken an existing model, extracted a free energy gradient from it, and then tested our principle on it.
+We begin with a generative model and a choice of non-equilibrium steady state, and then engineer a Markov-blanketted Langevin dynamics of the form (2.1.1) with prescribed $Q$ and $\Gamma$, showing in detail how these dynamics can be written as a flow on variational free energy.
+Since the periodically firing cells model above has been crafted in a bespoke fashion to have a prescribed $Q$, $\Gamma$, and non-equilibrium steady state, it would be interesting to contrast this with a given set of equations, validated by independent modelling work, and treat it using the above reasoning.
+That is to say, the example was explicitly designed so that the solenoidal term $Q$ and dissipative term $\Gamma$ are convenient (and, for simplicity, taken to be constant), making the inferential structure and thermodynamic quantities analytically transparent.
+This means that the example is not an independent empirical test of the FEP, but a carefully crafted illustration of what the formalism looks like when the required decomposition is assumed to exist.
+However—this top-down approach illustrates how our FEP might be used to model a system and explain some observed behaviours inferentially, once a dynamical system built around those behaviours is given.
+The examples in §3.1–3.2 are explicit constructions that exploit this fact: starting from a chosen non-equilibrium steady state and generative model, they show how the FEP provides a tractable representation of self-organisation (to a fixed point or limit cycle) as approximate Bayesian inference.
+A more demanding, ‘bottom-up’ demonstration would begin from an independently specified stochastic dynamical model—for example, a physically motivated Langevin equation—and then: (i) derive its steady-state density $p^{*}(x)$ and probability current, (ii) determine whether there exist fields $Q(x)$ and $\Gamma(x)$ satisfying the constraints of (2.1.1), and if so, (iii) read off the corresponding free energy functional and variational interpretation.
+This is precisely the kind of agnostic derivation that would address worries about circularity: instead of constructing a system to fit a free-energy-minimising flow, one would show that a pre-existing model can be recast in this way under clearly stated assumptions.
+Our central contribution in the present paper is therefore more modest but, we would argue, still substantive.
+Namely, we show that, whenever a decomposition such as (2.1.1) and a Markov blanket exist, one can construct an associated family of generative models and variational densities such that the physical dynamics can be written as descending a variational free energy functional.
+On this basis, the two toy systems are offered as worked examples of this constructive use of the FEP, rather than as empirical validations of its universality.
 4 Some preliminary philosophical considerations
@@
 This means that the FEP is what it says on the tin: it is a principle.
-Like the principles of least action or of maximum entropy, the FEP is a machinery that produces models [3].
+Like the principles of least action or of maximum entropy, the FEP is a piece of machinery that produces models [ramstead2022bayesian].
 Throughout, ‘as if’ marks an interpretive equivalence: internal flows that minimise variational free energy can be read as variational inference without presupposing explicit, propositional models inside the system.
@@
 If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed.
-As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
+As noted above, the technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
 If we now move to a meta-theoretical perspective and consider one particle (e.g., a scientist or philosopher) observing another particle, we have the interesting situation where the modeller may impute a particular generative model that best explains the mechanics of the observed particle.
 Is this meta-model a map or a territory?
-[^4] This meta-theoretical move is commonplace in practical applications of the FEP and is sometimes referred to as meta-Bayesian inference or, more simply, observing the observer [36].
+[^4] This meta-theoretical move is commonplace in practical applications of the FEP and is sometimes referred to as meta-Bayesian inference or, more simply, observing the observer [daunizeau].
 Practically, it involves optimising the parameters of a generative model, such that under the FEP (the ideal Bayesian observer assumption) the observable behaviour of the particle is rendered the most likely.
-This application of the FEP is often described as computational phenotyping [37].
+This application of the FEP is often described as computational phenotyping [schwartenbeck].
 Using a generative model to capture the formal structure of the environment—and to distinguish it from the model that is carried by particles (the variational density)—can be read as a new take on the nouvelle AI idea that physical systems are their own best models; or as a new take on the good regulator theorem.
-For instance, in work on robotics following the tradition of embodied cognition [38, 39], practitioners have eschewed the construction of agents with internal representations of some environment.
+For instance, in work on robotics following the tradition of embodied cognition [beer1995computational, brooks1991intelligence], practitioners have eschewed the construction of agents with internal representations of some environment.
 In these systems, the physics and geometry of the situation were sufficient to endow these agents with the capacity to couple to an environment.
@@
 Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself.
-In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in.
+In addition, the application of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in.
 Even if one could, the internal states would be forever sequestered behind blanket states.
 Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
-Here both observed and observer are described at the mechanistic level by flows on free energy flows; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.
+Here both observed and observer are described at the mechanistic level by flows on free energy gradients; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.
 5 A map of that part of the territory that behaves as if it were a map
 Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled).
-The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [40, 41].
-Some have referenced this fallacy in relation to prior work on the FEP [42, 43].
+The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [cartwright1984laws, cartwright1989nature].
+Some have referenced this fallacy in relation to prior work on the FEP [Bruineberg2020, VanEs2020].
 In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled.
-(See [6, 44, 45, 3] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification.
+(See [Andrews2021, andrews2022making, kirchhoff2022literalist, ramstead2022bayesian] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification.
 The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific, FEP-theoretic model) with the “territory” (the real-world system itself).
-This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [6, 44].
+This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [Andrews2021, andrews2022making].
 By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
-We argue that the FEP can be understood, metaphorically, a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map.
+We argue that the FEP can be understood, metaphorically, as a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map.
 In other words, the FEP provides us with tools to understand the mathematically striking property of self-organising systems: that they look as if they infer, track, reflect, or represent the statistics of their environment.
@@
 [^5] Here, we are assuming representationalist accounts of scientific practice.
-For alternatives, see [6].
+For alternatives, see [Andrews2021].
 Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics?
@@
 [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well.
-For a defence of the claim that physical systems are quite literally in the business of inference, see [46, 47].
-Related to this, views of physical systems implementing computations has been articulated [48, 49], as well as views where systems self-organise by computing predictions [50, 51, 26, 23].
+For a defence of the claim that physical systems are quite literally in the business of inference, see [Kiefer2017literal, Kiefer2020psycho].
+Related to this, views of physical systems implementing computations has been articulated [horsman2014does, aaronson2005guest], as well as views where systems self-organise by computing predictions [still2012thermodynamics, perunov2016statistical, ueltzhoffer2021drive, seifert2019stochastic].
 What is at stake with the FEP is an ‘as if’ description.
@@
 We remark that we do not claim to invalidate the map-territory fallacy itself.
-That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [44]).
+That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [andrews2022making]).
 Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification.
@@
 Taking all of the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system.
-This echoes seminal work in philosophy by Wittgenstein [52].
+This echoes seminal work in philosophy by Wittgenstein [wittgenstein2013tractatus].
 In his famous Tractatus Logico-Philosophicus, Wittgenstein set out to delimit the domain of sensible propositions by determining the general form of a proposition, an utterance able to carry a truth value.
@@
 6 Conclusion
-We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle.
+We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications of the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle.
 We have further argued that this allegation itself constitutes a map-territory fallacy fallacy.